### v1

# Generative Large Language Models in Automated Fact-Checking: A Survey

Abstract
The dissemination of false information across online platforms poses a serious societal challenge, necessitating robust measures for information verification. While manual fact-checking efforts are still instrumental, the growing volume of false information requires automated methods. Large language models (LLMs) offer promising opportunities to assist fact-checkers, leveraging LLM’s extensive knowledge and robust reasoning capabilities. In this survey paper, we investigate the utilization of generative LLMs in the realm of fact-checking, illustrating various approaches that have been employed and techniques for prompting or fine-tuning LLMs. By providing an overview of existing approaches, this survey aims to improve the understanding of utilizing LLMs in fact-checking and to facilitate further progress in LLMs’ involvement in this process.
Generative Large Language Models in Automated Fact-Checking: A Survey
Ivan Vykopal1,2, Matúš Pikuliak2, Simon Ostermann3, Marián Šimko2 1 Faculty of Information Technology, Brno University of Technology, Brno, Czech Republic 2 Kempelen Institute of Intelligent Technologies, Bratislava, Slovakia {name.surname}@kinit.sk 3 German Research Center for Artificial Intelligence (DFKI), Saarbrücken, Germany simon.ostermann@dfki.de
1 Introduction
The contemporary digital era has brought various challenges, including the spread of false information. The prevalence of social media has made this problem more challenging and a serious societal concern. An effective strategy for combating false information involves fact-checking Vlachos and Riedel (2014), mostly done manually by fact-checkers. However, the number of fact-checkers is limited, and the spread of false information is increasing Aïmeur et al. (2023). Hence, there is an effort to develop an automated fact-checking process by employing digital tools and LLMs that would assist fact-checkers Nakov et al. (2021a).
LLMs utilize large-scale datasets, paired with billions of parameters, to model language nuances and patterns of natural language. Furthermore, generative LLMs represent a specialized subset of LLMs tailored for text generation. Their application in manual fact-checking has demonstrated promising results in enhancing the efficiency and accuracy of fact-checking. With this study, we provide the first comprehensive summarization of approaches and techniques that incorporate generative LLMs in the fact-checking process.
Existing surveys have explored the needs of fact-checkers and tasks that can be automated Nakov et al. (2021a), task definitions, traditional methods, or BERT-like architectures Thorne and Vlachos (2018); Zeng et al. (2021). Moreover, several authors have examined existing datasets for fact-checking Guo et al. (2022). Another survey aimed to combat false information using LLMs and explore opportunities and challenges that LLMs bring Chen and Shu (2023b). However, this study did not explore methods used in detail, which presents an opportunity to address the role of generative LLMs in fact-checking more closely.
Our main contribution is to provide an overview of the utilization of generative LLMs in automated fact-checking. We present an overview of 70 papers with relevant approaches and novel prompting techniques for researchers who want to delve into LLM-aided information verification. We identify four main tasks and various proposed strategies to tackle them. Furthermore, we discuss future challenges and possible directions for utilizing LLMs in the information verification process.
2 Methodology
2.1 Selection
To identify a set of relevant articles, we employed a manual search of papers from NLP conferences, including ACL, EACL, NAACL and EMNLP. Along with these known conferences, we utilized a keyword search using Google Scholar and ArXiv.
In addition, we extended the created set to include papers referenced in related works and those that cited any article from our retrieved list of relevant papers. Additionally, papers submitted to CheckThat! Shared tasks were also involved in our survey. Overall, we collected 70 papers that employ generative LLMs for fact-checking.
2.2 Scope of the Survey
The survey is intended for researchers with expertise in the domain and a fundamental understanding of methods used across the NLP field. Our objective is to present an overview of techniques employed for fact-checking using generative LLMs. While fact-checking has evolved to include various modalities, our study delves only into the domain of textual and tabular information. Unlike other surveys that describe fact-checking datasets in detail, our focus lies on approaches, techniques, and associated challenges leveraging generative LLMs.
3 Tasks and Approaches Employed in Fact-Checking
In this section, we delve into the particular tasks of automated fact-checking, describing its definitions and the methods employed when engaging generative LLMs in the fact-checking process.
3.1 Check-Worthy Claim Detection
The task of check-worthy claim detection concerns identifying claims that contain potentially misleading information and that should thus be checked and verified. Selecting those claims is challenging, and there is no clear definition of a check-worthy claim. However, claims that contain information that may be false or misunderstood and are potentially harmful to society are usually considered to be check-worthy claims.
Generally, this task is considered a binary classification, the most common approach when employing LLMs. However, a finer-grained classification can be used to determine the harmfulness of a claim Wang et al. (2023b). Additionally, several authors have investigated the generative capabilities of LLMs to generate claims, where LLMs generate a set of verifiable claims from a given text, or questions, which aims to produce questions for further verification Chen et al. (2022). The advantage of using generative abilities, specifically the above techniques is that LLMs are able to extract key information and decompose complex statements into easily verifiable parts or questions.
3.2 Previously Fact-Checked Claims Detection
[…]
In previously fact-checked claims detection, LLMs are usually used to rank the already retrieved fact-checks, e.g., retrieved using the similarity of text embeddings. One possible application is to classify pairwise relationships between claims and retrieved fact-checks as textual entailment. In addition, LLMs can be employed to filter out irrelevant fact-checks that have been retrieved using another ranking approach.
3.3 Evidence Retrieval
The objective of evidence retrieval is to gather crucial information from trusted sources to assess the claim’s veracity, which can have multiple forms, e.g., text or tables. The purpose of evidence is to provide the knowledge needed for fact-checking to assess the veracity of the claims that are being fact-checked.
When incorporating LLMs into the fact-checking, the evidence fills knowledge gaps in LLMs, keeping them updated with up-to-date information they lack. In our paper, we refer to this setup as open-book setting, which allows LLMs to access external knowledge, such as Wikipedia or open web Schlichtkrull et al. (2023). However, selecting the knowledge source is crucial, and Wikipedia and the open web have several weaknesses. Thus, scientific reports Bhatia et al. (2021) are a more credible choice. Conversely, some approaches rely only on internal LLM knowledge and are termed closed-book settings.
The most employed approach for evidence retrieval is ranking. Alternative strategies explored the generative abilities of LLMs for query generation, in which LLM generates a search query for web search, and rationale selection, in which the LLM is tasked to select relevant sentences out of the previously retrieved documents.
3.4 Fact Verification & Fake News Detection
Fact verification focuses on determining the truthfulness of a claim or text based on the previously retrieved evidence. On the other hand, the fake news detection tries to classify a given input without the evidence based solely on the provided input.
The most prominent approach employing generative LLMs involves classifying a claim into three categories (supports, refutes, or not enough information) based on the given evidence. However, it can also be considered a binary or more fine-grained classification problem, e.g., six Buchholz (2023) or seven categories Hoes et al. (2023a). In addition to classification, explanation generation is a commonly used technique when incorporating generative LLMs for fact verification, where the goal is to explain the LLM’s decision and clarify what is misleading on a given claim.
4 Taxonomy of Methods
Figure 1: A taxonomy of methods employed when incorporating generative LLMs in the fact-checking process.
Generative LLMs have been employed in fact-checking through several approaches, which we divided into three categories: Classification & Regression, Generation, and Synthetic Data Generation. These categories differ in the granularity of the LLM’s output. In Classification & Regression, LLMs categorize inputs into discrete labels or continuous values. In Generation, the output is continuous text based on the input data. For Synthetic Data Generation, LLMs generate parts or entire datasets for training, validating and testing fact-checking methods. The proposed taxonomy is illustrated in Figure 1 and a list of relevant papers and detailed information is in Tables 2, 3, 4 and 4.
Classification & Regression and Generation are further categorized into prompting, fine-tuning and augmentation with external knowledge. Prompting leverages instructions to guide responses without LLM’s adaption, fine-tuning adapts LLMs to specific tasks via training on target tasks and knowledge augmentation enhances the responses of LLMs by incorporating up-to-date information.
4.1 Classification & Regression
In classification, LLMs predict categorical labels, while in regression LLMs are instructed to provide continuous values as predicted scores on various scales (0-1, 0-100, etc.). However, regression is often reframed into a soft classification problem due to the absence of datasets with score targets. In soft classification, the authors leverage the produced scores to determine the final class using thresholds (commonly 50%). The benefit of soft classification is that the threshold for the classes can be optimized based on the identified properties of LLMs.
Since the objective of most fact-checking tasks is to classify a claim into a certain category, classification is the most commonly used method in fact-checking, especially in check-worthy claim detection (e.g., yes or no) and fact verification (e.g., false, half true, true). Nevertheless, its potential use is also in other tasks, such as detecting previously fact-checked claims as an entailment classification between a given statement and previously fact-checked claims. In contrast, regression is widely employed in previously fact-checked claims detection and evidence retrieval, where the identified claims or evidence are ranked using a predicted score. Furthermore, regression is leveraged to a small extent in fact verification, intending to predict the truthfulness score for a given claim.
4.1.1 Prompting
In the era of generative LLMs, prompting has gained prominence, which is a method of instructing the LLM to condition the response to achieve the desired result. This method is employed in many NLP tasks, and several authors investigated the capabilities of prompting in the fact-checking domain. In classification, prompts with a list of possible categories are commonly used, while LLMs are instructed to choose the most probable label. These prompting techniques are particularly beneficial when datasets are limited across various languages or domains, making traditional fine-tuning methods challenging to use. This is especially a concern for fact-checking datasets which are limited mostly by their size.
Zero-Shot Prompting.
In zero-shot prompting, LLMs are directly prompted with the instruction without providing examples. This is particularly beneficial when there is a lack of extensive training data and quality samples to serve as in-context examples. However, zero-shot prompting imposes several limitations. One of them is that instructions need to be clear and describe the task context to perform it correctly. It is crucial to introduce all the terminology within the prompt to mitigate ambiguities, e.g., attributes of check-worthy claims. Due to these challenges, prompt engineering is an essential part of experiments that need to be conducted, especially for fact-checking.
Another limitation is that the instruction wording and style can significantly affect LLM performance. To address this issue, various prompt styles have been proposed. In fact-checking, especially role specification Li and Zhai (2023) and JSON format Ma et al. (2023) have shown promise in enhancing performance. Role specification involves instructing LLM to act based on a defined role, such as fact-checker, while in the JSON format, the instruction adheres to JSON structure (e.g., Check the ${claim}\nAnswer:$). Role specification aids LLMs in better understanding the task, while the JSON facilitates easier post-processing due to its structured response format.
Since fact-checking poses a challenging task for zero-shot prompting and provided answers by LLMs may not always be based on facts, Zeng and Gao (2023) utilized multiple prompt variants to assess the veracity label of each claim. These prompt variants differ in only one word and represent various relationships between the claim and whether it is true, unclear or false, to avoid ambiguities in the output of the LLM. This technique verifies the consistency of LLMs’ responses to a given claim.
Few-Shot Prompting.
Zero-shot prompting has one major limitation: the model is not always able to capture the essence of the task based only on the given instruction. Therefore, few-shot prompting is often used to enhance LLM performance by providing demonstrations necessary for complex tasks like fact-checking. LLMs efficiently learn the nuances of the task during inference utilizing provided labelled examples. Applying few-shot prompting is similar to zero-shot, whereas the authors also investigated the number of samples provided to the LLM Cao et al. (2023); Zeng and Gao (2023, 2024).
Chain-of-Thought.
The Chain-of-Thought (CoT) technique harnesses the power of LLMs to provide logical reasoning steps. This reasoning is performed before the actual class is provided, with the LLM exploiting its reasoning capabilities by incorporating intermediate reasoning steps for the final assessment. The CoT can be used in zero-shot Hu et al. (2023); Chen and Shu (2023a) and few-shot Liu et al. (2024a) settings, where authors include examples of the steps that reasoning should involve, e.g., follow-up questions Sawiński et al. (2023).
4.1.2 Fine-Tuning
Fine-tuning represents the method for adapting LLMs to a specific task or domain. It is mainly employed when the LLM does not have the necessary capabilities to address a given task using prompting, especially in cases where the LLM has no expertise in the selected domain or the task is complex. However, due to a large number of parameters, fine-tuning is usually only a concern for smaller LLMs. For larger LLMs, fine-tuning is often not feasible due to the limited resources, and also, it is not necessary, since these LLMs have improved capabilities for solving a wide range of tasks, even without further fine-tuning. In fact-checking, fine-tuning is primarily employed for ranking tasks, e.g., previously fact-checked claims detection and evidence retrieval. However, it is also used for other tasks beyond these applications.
LLM fine-tuning can be done in multiple settings and one of them is LLM fine-tuning on a single task. The advantage of such training is the specialization of the LLM to only a specific task and hence the improved ability of the LLM to address it. Conversely, the drawback of this approach lies in the fact that the LLM loses the ability to solve other tasks, even if they are from the same domain (e.g., fact-checking), known as catastrophic forgetting McCloskey and Cohen (1989). In this context, addressing multiple tasks necessitates fine-tuning a specific LLM for each task. One technique for LLM fine-tuning on a single task involves leveraging the logits of the particular token produced by the LLM, such as true token, which can be used for document ranking Pradeep et al. (2021). We refer to document ranking based on the probability of the true token as classification ranking because the LLM has been trained in a binary classification setting. Building upon classification ranking, multiple authors utilized additional steps Jiang et al. (2021); Pradeep et al. (2021); Fernández-Pichel et al. (2022). The classification ranking approach can be further employed to rank evidence in evidence retrieval and to extract pertinent sentences from retrieved evidence Pradeep et al. (2020).
Another option for fine-tuning is multitask fine-tuning, involving multiple tasks during training Du et al. (2022). This fine-tuning method aims to create a more generalizable model while exploiting knowledge transfer between tasks. This technique is not intended to improve the performance of a single task, but of all the tasks involved, where each task can positively affect the others. The other benefit compared to training on a single task is that the LLM is specialized for multiple tasks simultaneously and hence it is not necessary to fine-tune separate LLMs for each task.
LLMs can be fine-tuned with target labels leveraging various prompts. For fine-tuning encoder-decoder LLMs, authors mainly employed a prompt of the form "Claim: c Evidence: e Target:" Sarrouti et al. (2021), while for fine-tuning decoder-only LLMs, the claim alone is often utilized as input Sawiński et al. (2023). Since current LLMs are usually instruction-tuned, including the instruction along with the statement is common.
4.1.3 Augmentation with External Knowledge
Nevertheless, multiple weaknesses persist in prompting techniques, including a frequent lack of access to up-to-date information. Up-to-date information is crucial for accurate fact-checking, and its absence can lead to inaccurate responses. To counter this issue, the evidence is commonly incorporated into the prompt. Since multiple pieces of evidence may be available to verify a statement’s veracity, there are two schemes that combine multiple pieces of evidence Jiang et al. (2021). The first method verifies the claim sentence by sentence, and aggregates predicted labels for the final prediction. In the second approach, all relevant sentences are employed within a single query, surpassing the performance of the first approach. Unlike textual information, evidence can also be in the form of tabular data, which is often linearized before providing it to LLMs Zhang et al. (2024b).
In contrast to incorporating retrieved evidence from the evidence retrieval task, it is common to employ methods for fact-checking, where LLMs interact with external tools, such as search engines or external knowledge bases Zhang and Gao (2023); Yao et al. (2023); Quelle and Bovet (2023). These techniques provide several benefits, the most significant being that the LLM always have access to the most up-to-date information, which is crucial for fact-checking.
4.2 Generation
LLMs exhibit impressive generative abilities, aimed at producing meaningful continuous text. This generation capability is beneficial for various NLP tasks, including fact-checking. Moreover, the outputs in natural languages enable the integration of LLMs into a wide range of applications, streamlining the fact-checking process. Additional advantages include the ability of the LLM to handle diverse topics and to adapt to different contexts, which is particularly valuable for fact-checking where there is a high diversity of topics.
Generation has the potential to be used across all fact-checking tasks, but especially in the fact verification, to explain misleading information within the claim.
4.2.1 Prompting
Prompting in generation involves providing the instruction to guide the LLM in producing meaningful coherent text. Prompting techniques are commonly used for generation, due to the absence of labelled data for LLM training, which are challenging to obtain. Furthermore, LLMs with billions of parameters are capable of addressing even complex tasks that are part of fact-checking without further fine-tuning.
Zero-Shot Prompting.
The same issues arise for the use of zero-shot prompting in generation as were identified for classification and regression, specifically that the style and content of the instruction affect the expected output. Since analyzed claims may be challenging for fact verification, LLMs can be instructed to reformulate a given claim into a neutral style that is used for the final assessment Wu and Hooi (2023). This approach enhances LLMs’ resistance to adversarial attacks.
In generation, both the input and the format and order of the output affect the LLM’s performance. In fact-checking, where the LLM is instructed to provide an explanation for its assessment, whether the explanation is generated before or after the final assessment is crucial. If the final assessment is produced before the explanation, the LLM may focus on matching the final score or category, resulting in an inaccurate explanation. Conversely, generating the explanation first allows it to reason behind the LLM’s answer, leading to a more accurate and coherent assessment.
In zero-shot prompting for evidence retrieval, LLMs can be instructed to generate search queries for search engines to retrieve relevant results. For instance, Prieto-Chavana et al. (2023) utilized prefixes, e.g., Search query:, to guide LLMs in producing queries. This approach allows LLMs to extract essential parts of a claim for better retrieval, avoiding the drawbacks of using the entire claim, which can result in irrelevant evidence. Another approach for evidence retrieval is rationale selection, where LLMs aim to select relevant parts of the retrieved evidence or their own knowledge, such as selecting relevant sentences Evans et al. (2023); Tan et al. (2023). Simultaneously, these two approaches complement each other, whereby we can obtain evidence from search engines using generated queries and reduce the evidence to only relevant parts using rationale selection.
Few-Shot Prompting.
Few-shot prompting is often used for simpler sub-tasks, such as query or claim generation. This approach extends to other tasks, e.g., explanation generation in fact verification. However, achieving better performance in these tasks requires quality examples. By using a few demonstrations, the LLM can capture how to accurately extract verifiable claims from a given input, model the desired structure of the answer, or recognize the features that generated questions should have. Moreover, in-context examples are commonly employed to improve the LLM’s understanding of the task.
Possible methods of using in-context examples in prompting are span detection Gangi Reddy et al. (2022), aiming to identify the exact boundaries of the check-worthy claims within the text, and decomposition of input text into sub-claims Kamoi et al. (2023b); Li et al. (2023a). A few demonstrations can assist LLM in comprehending what characteristics a check-worthy claim is supposed to have, and what its variability might be. With this approach, it is important to ensure that the examples capture a wide range of characteristics of check-worthy claims, to eliminate the false positive rate. The advantage of span detection lies in the fact that the LLM is instructed to select the exact wording of the claim from the input. However, when decomposing the input into sub-claims, the LLM may reformulate the claim potentially altering its meaning and truth value.
Additionally, few-shot prompting is not limited to only check-worthy claim detection, but is also employed for evidence retrieval, for generating search queries Li et al. (2023a) or rationale selection Wang et al. (2023a). While, the mentioned advantages, disadvantages and attributes of few-shot prompting are relevant for all fact-checking tasks.
Chain-of-Thought.
CoT represents reasoning steps to achieve the desired answer. Similarly to classification and regression, CoT can be used in a generation without an example or with several demonstrations. However, CoT is not widely employed in generation, since the aim of generation is already to provide meaningful text and explanation is often part of it.
Few-shot CoT can, for example, be used in a check-worthy claim detection task in producing normalized claim based on a given social media post Sundriyal et al. (2023). However, the drawback of a single normalized claim arises from the potential presence of multiple facts within these posts, necessitating verification to accurately asses their veracity.
4.2.2 Fine-Tuning
Fine-tuning is not as widespread in generation as in classification or regression, since obtaining quality data for this purpose is challenging. However, fine-tuning can be used if there is a similar task in another domain with more available data to train LLMs. For example, summarization can be leveraged as part of explanation generation in fact verification Russo et al. (2023). The flaw of summarization for fact-checking is that this approach is unable to capture the essential parts of the retrieved evidence that are necessary to explain what is misleading in the claim and the summaries generated are not as convincing as actual fact-checks.
In addition, fine-tuning was employed in the generation of sub-questions to verify a given claim Chen et al. (2022) and for next-hop prediction Malon (2021), which are both necessary for predicting the veracity label correctly. On the one hand, the generation of sub-questions aims to verify the truth of a claim it is necessary to ask the proper questions. On the other hand, next-hop prediction is crucial, when retired evidence introduces entities requiring further elaboration.
4.2.3 Augmentation with External Knowledge
Researchers have explored various methods to mitigate hallucinations and outdated information in fact-checking. One effective method is Retrieval Augmented Generation (RAG), which integrates information retrieval with the LLM, leveraging external sources to enhance accuracy. This technique proves powerful in combating hallucinations by using credible information extracted from sources like scientific repositories Leippold et al. (2024). Moreover, RAG can be fine-tuned by jointly training the retriever and LLM, where the LLM provides supervisory signals for training the retriever Izacard et al. (2023); Zeng and Gao (2024).
4.3 Synthetic Data Generation
Synthetic data generation aims to generate an entire dataset or its part with LLMs. Unlike generation for specific tasks, synthetic data generation precedes the fact-checking pipeline and is not directly involved in fact-checking tasks. This approach is advantageous when there are no datasets for some tasks or languages or if there is a limited number of such data. In fact-checking, it can be used for generating datasets with disinformation Chen and Shu (2023a); Vykopal et al. (2024) or fake-news Jiang et al. (2023); Huang and Sun (2023).
Additionally, synthetic data generation can augment datasets by filling gaps in sparse categories or enhancing variability in the data. For instance, LLM can generate debunked claims to extend datasets Singh et al. (2023) or create counterparts for claim-evidence pairs with contrasting veracity but high word overlap Zhang et al. (2024a).
5 LLM Pipelines
It is common that LLMs are used in more complex orchestrated pipelines where the input is being processed in multiple steps and usually the output of one step is used as the input for another. Several authors proposed pipelines consisting of multiple fact-checking tasks and employed multiple described methods to provide complex answers necessary for fact-checkers.
One common approach used by multiple authors consists of decomposing the given claim into easily verifiable sub-claims and predicting the veracity of each generated sub-claim. Zhang and Gao (2023) employed the LLM to verify each sub-claim using a generated set of questions and answers. In this case, the LLM can utilize its internal knowledge or retrieve information from the web to answer these questions, assessing its confidence level before responding. Additionally, Wang and Shu (2023) instructed the LLM to define predicates and follow-up questions. Once the answers are provided, the LLM utilizes a context comprising predicted follow-up questions and their answers to assess the veracity of each predicate, which are further used for the final prediction and explanation.
In addition, the decomposition of the claim into sub-claims can be omitted and the LLM can be directly instructed to generate questions and answers for veracity prediction. LLMs can be leveraged to generate questions along with answers, veracity prediction for each answer, and overall veracity label for a given claim Chakraborty et al. (2023). In this case, the LLM decided whether it supports or refutes the claim or there is not enough information. Moreover, claims can be decomposed into binary questions (yes/no) with corresponding answers, while these answers are used as input for generating interpretable logic clauses to debunk misinformation Liu et al. (2024a).
Besides standard generative LLMs, Pan et al. (2023b) explored a programming LLM, leveraging its program generation capabilities to create a verification program. This approach decomposes the input claim into questions using one generative LLM and then employs programming LLM to generate a reasoning program to guide verification. Subsequently, the generated code is executed using the corresponding sub-task function, and the final veracity prediction is produced.
Since LLMs tend to produce hallucinated and unfaithful explanations, generative LLMs can also be used to enhance faithfulness refinement Kim et al. (2024). This approach harnesses two LLM debaters, which iteratively search for errors and flaws in the explanation. Subsequently, all identified errors and proposed corrections from both debaters are used to adjust the final justification. This process is valuable in fact-checking, which requires a convincing and plausible explanation, as a human fact-checker normally would do.
6 Multilinguality
False information is not only a problem in English-speaking countries but a worldwide phenomenon, applying LLMs across multiple languages is necessary. However, most existing work has focused on applying LLMs to only a single language, especially English.
In evaluating the performance of LLMs across different languages, most researchers adopted approaches similar to those used in single-language assessment, using the techniques described in Section 4. Specifically, when evaluating the fact-checking capability of generative LLMs in various languages, the authors predominantly utilized instructions in a single language, typically English, with the claims presented in their original languages Du et al. (2022); Agrestia et al. (2022); Hyben et al. (2023); Cao et al. (2023); Huang and Sun (2023); Pelrine et al. (2023); Li and Zhai (2023). This methodology allowed for the use of the same prompts without adapting it to the language of the analyzed claim. Moreover, the instruction can be in the same language as the analyzed claim, which has been observed to enhance the accuracy of LLMs in identifying false information Hu et al. (2023). However, this approach has its drawbacks. It requires predefined instructions for each supported language, and it also relies on accurate language detectors.
Another method employed in the case of multilingualism and the use of generative LLMs, focuses on fine-tuning LLMs on multilingual data, to create a generalized LLM capable of processing inputs in multiple languages Quelle and Bovet (2023). Additionally, LLMs can be trained in one language, particularly a high-resource one, and then evaluated in other languages, a process known as cross-lingual transfer Pikuliak et al. (2021).
7 Future Directions and Challenges
Knowledge-Augmented Strategies.
Techniques like ReAct Yao et al. (2023) and RAG Lewis et al. (2021), which incorporate external knowledge or tools, have shown promise in addressing complex NLP tasks, including fact-checking. Despite their potential, these methods are still underexplored in fact-checking, only 5 out of 70 surveyed papers. Future directions lie in harnessing these techniques to enhance the efficiency and accuracy of fact-checking, which brings opportunities for detecting false information.
Multilingual Fact-Checking Benchmarks.
Evaluating LLMs in a multilingual context is necessary for fact-checking since false information is a global issue. Developing fact-checking benchmarks in multiple languages can help to assess and improve the LLM’s cross-lingual capabilities. Since new LLMs emerge rapidly, the established benchmarks will aid in faster and more efficient evaluation of both current and future LLMs, thus empowering fact-checkers to always use state-of-the-art models.
Fact-Checking in Real Time.
In the current research landscape, most approaches act reactively, often after a claim gains traction on social media platforms. Leveraging generative LLMs can enhance this process by enabling continuous monitoring and instant analysis of emerging information across various platforms. To ensure LLMs keep up-to-date and effective in fact-checking, they need to be integrated with tools allowing LLMs to access the newest information in real-time. This proactive approach can identify and flag false claims and enhance the accuracy and reliability of information dissemination.
Interactive Fact-Checking.
Interactive fact-checking tools, especially those exploiting generative LLMs, represent a promising future direction in combating false information. Users can engage in dialogue with an LLM to explore the accuracy of claims, ask follow-up questions or request explanations of its decisions. The development of interactive fact-checking tools offers an excellent opportunity to empower fact-checkers in effective false information mitigation.
8 Conclusion
The rapid development of generative LLMs has stimulated considerable interest in their potential utilization in fact-checking. Through a systematic review of 70 research papers, this study gave an overview of methods that incorporate LLMs into information verification. This survey discusses the techniques and approaches explored in the reviewed literature, which not only provides a comprehensive overview of current methodologies but also presents the basis for future research efforts to improve existing methods and explore new frontiers in LLM-assisted fact-checking.
Limitations
Focus on Generative LLMs.
This survey paper aims to provide an overview of the approaches and techniques employed in fact-checking while covering only generative LLMs. There are various types of language models, but in recent years, the majority of the research in the NLP domain has focused on LLMs, especially on generative LLMs.
Textual Information Only.
Although false information is not only an issue in textual form, but it can involve various modalities, such as images, videos, speech, etc., in our survey we focus only on the problem of verifying textual information. Therefore, we excluded all papers that did not involve text processing in their approach.
Acknowledgements
This research was partially supported by DisAI - Improving scientific excellence and creativity in combating disinformation with artificial intelligence and language technologies, a project funded by Horizon Europe under GA No.101079164, by the Central European Digital Media Observatory 2.0 (CEDMO 2.0), a project funded by the European Union under the Contract No. 101158609, and by the MIMEDIS, a project funded by the Slovak Research and Development Agency under GA No. APVV-21-0114.
Appendix A Datasets & Languages
After the rise of social media, the necessity for the development of datasets aimed at verifying information began to emerge. The first existing datasets started to appear in early 2010 and their availability increased over time. Many of these datasets have been involved in experiments using LLMs in fact-checking. Several of them come from Conference and Labs of the Evaluation Forum (CLEF) and specifically CheckThat! Lab focused on various tasks related to information verification, political bias of news articles, and anti-social behaviour, such as hate speech or propaganda detection.
In addition, multiple datasets focused thematically on different topics, e.g. COVID-19, the climate crisis, or political claims. Several datasets were developed from data collected from fact-checking organizations, for example, MultiClaim Pikuliak et al. (2023b) or LIAR Wang (2017). Furthermore, multiple authors engaged their own data collection resources, such as PolitiFact [^1] https://www.politifact.com/ or Google Fact Check Tools [^2] https://toolbox.google.com/factcheck/explorer, which collects verified claims from the Internet based on the ClaimReview schema [^3] https://schema.org/ClaimReview.
A key attribute for identifying false claims in cross-lingual settings is the existence of multilingual datasets, comprising multiple languages or a combination of language-specific datasets. Previously, most datasets have focused on one or a few languages, specifically, English and Arabic are the most represented languages. Recently, there have been efforts to create multilingual datasets that include a broader range of languages, including e.g. MultiClaim Pikuliak et al. (2023b), X-Fact Gupta and Srikumar (2021) or Pinocchio Hu et al. (2023). A list of all identified datasets employed for the LLM experiments is shown in Table LABEL:tab:datasets.
Concerning linguistic scope, the vast majority of papers are concentrated on English ($57\times$57 ×) or a single language like Chinese ($1\times$1 ×) or Arabic ($1\times$1 ×). Out of the total papers surveyed, 12 analyzed more than one language, predominantly combining English with another language. Notable, only four investigated the employing LLMs across ten or more languages. Nevertheless, the most extensive language coverage in LLM experiments involved 114 languages Setty (2024), or 46 languages in the fact verification task.
Appendix B Fact-Checking Benchmarks
Several benchmarks have been developed to assess the performance of LLMs in fact-checking tasks. Hu et al. (2023) introduced the Pinocchio benchmark, encompassing seven domains related to factual knowledge and reasoning, evaluating ten LLMs with different prompting strategies. In contrast, Bang et al. (2023) focused on assessing ChatGPT’s reasoning, hallucination, and interactivity capabilities across eight tasks, including misinformation detection. Another benchmark, TrustLLM, scrutinized the trustworthiness and safety features of 30 LLMs Sun et al. (2024).
Appendix C Model Analysis
Various researchers have focused on different types of LLMs within the reviewed papers. In contrast, these could be divided into two main categories: Encoder-Decoder (Seq2Seq) and Decoder-only (as shown in Figure 2). Although the first Seq2Seq models already emerged in 2019, their first applications in fact-checking were seen in 2021, when the problem of automated fact-checking began to receive increasing attention. The most commonly used Seq2Seq models include the T5 model, while we also encounter the T0, Flan-T5, or multilingual mT5.
Figure 2: The hierarchy of identified models in surveyed papers.
In contrast, the second category includes Decoder-only LLMs, which we further divided into GPT-base, LLaMA-based and Other models, similarly to Minaee et al. (2024).
  1. 1.
    The GPT Family. Generative Pre-trained Transformers (GPT) is a family of LLMs developed by OpenAI and includes mostly closed-source models, accessible by paid API. This category includes all GPT models (e.g. GPT-3, GPT-3.5, or GPT-4), CODEX. In addition to these closed-source models, there are also several open-sourced models, such as the GPT-Neo model.
  2. 2.
    The LLaMA Family. The LLaMA family is another collection of LLMs based on the LLaMA or Llama2 foundation models developed by Meta. This category also includes other models derived from LLaMA models, such as Alpaca, Vicuna, or even Mistral.
  3. 3.
    Other models. The last category consists of models that do not fall under the GPT or LLaMA family. Within this category, various models have gained the attention of researchers. Examples include the multilingual BLOOM or the OPT, PaLM, and PaLM2 models. Along with these models, several authors also investigated Atlas, a retrieval-augmented language model.
Figure 3 depicts LLMs employed in more than five papers, accompanied by the respective count of papers each model utilized. A total of 35 various LLMs were employed in the collection of 70 papers.
Figure 3: The occurrence frequency of the models in the reviewed papers. We only present models used in more than five papers.
Table 1: A list of datasets utilized in experiments with LLMs for fact-checking.
Name     Check-worthy claim detection     Previously fact-checked claims detection     Evidence retrieval     Fact verification # Lang. Size Citation
ArCOV19-Rumors 1 138 Haouari et al. (2021)
ArFactEx 1 100 Althabiti et al. (2024)
AveriTec 1 5k Schlichtkrull et al. (2023)
BoolQ-FV / AmbiFC 1 11k Glockner et al. (2023)
CHEF 1 10k Hu et al. (2022)
CLAN 1 6k Sundriyal et al. (2023)
CLEF-2020 3 13K Barrón-Cedeno et al. (2020)
CLEF-2021 5 22k Nakov et al. (2021b)
CLEF-2022 7 36k Nakov et al. (2022)
CLEF-2023 3 63k Alam et al. (2023)
CLIMATE-FEVER 1 2k Diggelmann et al. (2021)
COVID-19 Scientific 1 142 Lee et al. (2020)
COVID-Fact 1 4k Saakyan et al. (2021)
ChatGPT-FC 1 22k Li et al. (2023b)
Check-COVID 1 2k Wang et al. (2023a)
ClaimBuster 1 24k Arslan et al. (2020)
ClaimDecomp 1 1k Chen et al. (2022)
Climate Feedback 1 N/A -
CoAID 1 5k Cui and Lee (2020)
Constraint 1 11k Patwa et al. (2021)
Data Common X N/A -
e-FEVER 1 68k Stammbach and Ash (2020)
EX-FEVER 1 61k Ma et al. (2024)
EnvClaims 1 29k Stammbach et al. (2023)
ExClaim 1 4k Gurrapu et al. (2022)
FEVER 1 185k Thorne et al. (2018)
FEVEROUS 1 87k Aly et al. (2021)
FM2 1 13k Eisenschlos et al. (2021)
FactEX 1 12k Althabiti et al. (2023)
FakeNewsNet 1 23k Shu et al. (2019)
FullFact 1 N/A -
GermEval 2021 1 1k Risch et al. (2021)
Global-LIAR 1 600 Mirza et al. (2024)
HealthVer 1 14k Sarrouti et al. (2021)
HoVer 1 26k Jiang et al. (2020)
LESA-2021 1 379k Gupta et al. (2021)
LIAR 1 13k Wang (2017)
LIAR++ 1 6k Russo et al. (2023)
LIAR-NEW 2 2k Pelrine et al. (2023)
LLMFake 1 100 Chen and Shu (2023a)
Labeled Unreliable News 1 74k Rashkin et al. (2017)
Monant Medical Misinformation Dataset 1 51k Srba et al. (2022)
MS-MARCO 1 1011k Bajaj et al. (2018)
MSVEC 1 200 Evans et al. (2023)
MultiClaim 27/39 28k Pikuliak et al. (2023a)
NewsClaims 1 889 Gangi Reddy et al. (2022)
Pinocchio 27 21k Hu et al. (2023)
PolitiFact 1 N/A -
PolitiHop 1 500 Ostrowski et al. (2021)
PubHealth 1 12k Kotonya and Toni (2020)
RAWFC 1 2k Yang et al. (2022)
SciFact 1 1k Wadden et al. (2020)
Snopes 1 N/A -
Symmetric 1 1k Schuster et al. (2019)
TREC 2020 Health Misinfo. 1 46 Clarke et al. (2020)
TREC-COVID 1 171k Roberts et al. (2021)
TabFact 1 18k Chen et al. (2020)
Twitter15 1 2k Liu et al. (2015)
Twitter16 1 778 Ma et al. (2016)
VitaminC 1 489k Schuster et al. (2021)
WatClaimCheck 1 34k Khan et al. (2022)
Weibo21 1 9k Nan et al. (2021)
WiCE 1 7k Kamoi et al. (2023a)
X-Fact 26 31k Gupta and Srikumar (2021)
Paper Approaches Techniques # Lang.
Gangi Reddy et al. (2022) - classification - span detection - fine-tuning - zero-shot - few-shot 1
Agrestia et al. (2022) - binary classification - multilabel classification - fine-tuning 4
Du et al. (2022) - binary classification - fine-tuning 6
Chen et al. (2022) - claim decomposition - fine-tuning 1
Sawiński et al. (2023) - binary classification - fine-tuning - zero-shot - few-shot - few-shot CoT 1
Kamoi et al. (2023b) - text generation - few-shot 1
Li et al. (2023a) - text generation - few-shot 1
Hyben et al. (2023) - binary classification - fine-tuning - zero-shot - few-shot 10
Cao et al. (2023) - binary classification - zero-shot - few-shot - zero-shot CoT 2
Wang et al. (2023b) - 3-class classification - zero-shot 1
Sundriyal et al. (2023) - text generation - few-shot - zero-shot CoT 1
Setty (2024) - binary classification - zero-shot 114
Table 2: Work addressing the task of check-worthy claim detection in the context of fact-checking. Each paper includes the utilized approaches, techniques and the number of languages analyzed.
Paper Approaches Techniques # Lang.
Shliselberg et al. (2022) - ranking task - fine-tuning 1
Singh et al. (2023) - ranking task - fine-tuning - zero-shot 1
Choi and Ferrara (2023) - 3-class classification - zero-shot - zero-shot CoT - few-shot CoT 1
Neumann et al. (2023) - ranking task - fine-tuning 1
Choi and Ferrara (2024) - 3-class classification - fine-tuning - zero-shot 1
Table 3: A list of papers addressing the task of detecting previously fact-checked claims. We include approaches, techniques and a number of languages for each paper.
Paper Approaches Techniques Knowledge base # Lang.
Pradeep et al. (2020) - ranking - rationale selection - fine-tuning - MS-MARCO 1
Pradeep et al. (2021) - classification ranking - fine-tuning - given corpus 1
Sarrouti et al. (2021) - classification ranking - fine-tuning - search engine 1
Jiang et al. (2021) - classification ranking - fine-tuning - search engine 1
Fernández-Pichel et al. (2022) - classification ranking - fine-tuning - Med-MARCO 1
Kamoi et al. (2023b) - rationale selection - fine-tuning - Wikipedia 1
Prieto-Chavana et al. (2023) - query generation - fine-tuning - zero-shot - few-shot - search engine 1
Schlichtkrull et al. (2023) - question generation - few-shot - search engine 1
Wang et al. (2023a) - rationale selection - few-shot - scientific papers 1
Evans et al. (2023) - rationale selection - zero-shot - scientific papers 1
Tan et al. (2023) - rationale selection - few-shot - search engine 1
Pan et al. (2023a) - question generation - few-shot - Wikipedia - LLM knowledge 1
Setty (2024) - question generation - zero-shot - search engine - fact-checks 114
Table 4: A list of relevant works focused on evidence retrieval. For each analyzed work we also present the knowledge base employed.
{longtabu}
to lllll Paper Approaches Techniques Used Evidence # Lang.
Pradeep et al. (2020) - 3-class classification - fine-tuning - open-book 1
Petroni et al. (2021) - 3-class classification - fine-tuning - open-book 1
Pradeep et al. (2021) - 3-class classification - fine-tuning - open-book
Sarrouti et al. (2021) - 3-class classification - fine-tuning - open-book 1
Malon (2021) - 3-class classification - next hop prediction - fine-tuning - open-book 1
Jiang et al. (2021) - multi-class classification - fine-tuning - open-book 1
Tran and Kruschwitz (2022) - 4-class classification - fine-tuning - open-book 2
Bhatia et al. (2021) - binary classification - 3-class classification - explanation generation - fine-tuning - open-book - closed-book 1
Hoes et al. (2023a) - binary classification - 6-class classification - zero-shot - closed-book 1
Kamoi et al. (2023b) - binary classification - fine-tuning - few-shot - open-book 1
Schlichtkrull et al. (2023) - 3-class classification - explanation generation - few-shot - open-book - closed-book 1
Buchholz (2023) - 6-class classification - fine-tuning - zero-shot - closed-book 1
Wang et al. (2023a) - 3-class classification - few-shot - open-book 1
Pan et al. (2023b) - binary classification - text generation - explanation generation - zero-shot - few-shot - zero-shot CoT - few-shot CoT - few-shot self-ask - gold evidence - open-book - closed-book 1
Izacard et al. (2023) - 3-class classification - fine-tuning (RAG) - open-book 1
Huang and Sun (2023) - binary classification - 3-class classification - explanation generation - zero-shot - zero-shot CoT - self-aware prompt - closed-book 2
Jiang et al. (2023) - binary classification - regression - explanation generation - zero-shot - closed-book 1
Chen and Shu (2023a) - binary classification - zero-shot - zero-shot CoT - closed-book 1
Quelle and Bovet (2023) - 4-class classification - 7-class classification - ReAct - open-book - closed-book 14
Ma et al. (2023) - 3-class classification - explanation generation - zero-shot - few-shot - open-book 1
Guan et al. (2023) - regression - fine-tuning - zero-shot - open-book - closed-book 1
Hu et al. (2023) - 3-class classification - zero-shot - few-shot - zero-shot CoT - few-shot CoT - self-refinement - self-consistency - open-book 27
Wu and Hooi (2023) - binary classification - zero-shot - closed-book 1
Evans et al. (2023) - 3-class classification - regression - zero-shot - open-book 1
Althabiti et al. (2023) - explanation generation - fine-tuning - closed-book 1
Chakraborty et al. (2023) - binary classification - 3-class classification - explanation generation - zero-shot - closed-book 1
Bang et al. (2023) - explanation generation - zero-shot - closed-book 1
Tan et al. (2023) - 3-class classification - explanation generation - zero-shot - open-book 1
Zeng and Gao (2023) - 3-class classification - fine-tuning - zero-shot - few-shot - self-consistency CoT - open-book 1
Wang and Shu (2023) - 3-class classification - text generation - few-shot - CoT - self-ask - FOLK - open-book 1
Pelrine et al. (2023) - binary classification - regression - explanation generation - fine-tuning - zero-shot - few-shot - open-book - closed-book 3
Zhang and Gao (2023) - 3-class classification - 6-class classification - text generation - few-shot - few-shot CoT - search-augmented CoT - ReAct - HiSS - open-book - closed-book 1
Li et al. (2023b) - regression - explanation generation - zero-shot - closed-book 1
Hoes et al. (2023b) - binary classification - 4-class classification - zero-shot - closed-book 1
Pan et al. (2023a) - binary classification - explanation generation - zero-shot - few-shot - few-shot CoT - open-book - closed-book 1
Russo et al. (2023) - explanation generation - fine-tuning - open-book 1
Li and Zhai (2023) - 6-class classification - zero-shot - closed-book 2
Vergho et al. (2024) - binary classification - regression - explanation generation - zero-shot - closed-book 2
Zeng and Gao (2024) - 3-class classification - explanation generation - fine-tuning (RAG) - few-shot - open-book 1
Leippold et al. (2024) - binary classification - 5-class classification - 7-class classification - 12-class classification - RAG - open-book - closed-book 1
Kao and Yen (2024) - explanation generation - fine-tuning - closed-book 1
Zhang et al. (2024a) - 3-class classification - few-shot - gold evidence - open-book 1
Althabiti et al. (2024) - 3-class classification - explanation generation - zero-shot - open-book 1
Zeng and Zubiaga (2024) - 3-class classification - fine-tuning - few-shot - open-book 1
Mirza et al. (2024) - binary classification - 3-class classification - zero-shot - closed-book 1
Sun et al. (2024) - 3-class classification - zero-shot - open-book 1
Zhang et al. (2024b) - binary classification - fine-tuning - zero-shot - few-shot - open-book 1
Cekinel and Karagoz (2024) - explanation generation - fine-tuning - open-book 1
Liu et al. (2024b) - binary classification - explanation generation - zero-shot - closed-book 1
Kim et al. (2024) - binary classification - explanation generation - MADR - zero-shot - CoT - self-refine - open-book 1
Sanyal et al. (2024) - binary classification - zero-shot - open-book 1
Setty (2024) - regression - explanation generation - zero-shot - open-book 46
Liu et al. (2024a) - binary classification - 6-class classification - zero-shot - few-shot - zero-shot CoT - few-shot CoT - TELLER - open-book - closed-book 1
A comprehensive list of works for the fact verification task. Most of the surveyed papers address this task. In the table, we provide used evidence as open and closed-book, based on either the use of an external source or the knowledge embedded in the parameters.

### v2

@@
 Abstract
-The dissemination of false information across online platforms poses a serious societal challenge, necessitating robust measures for information verification.
-While manual fact-checking efforts are still instrumental, the growing volume of false information requires automated methods.
-Large language models (LLMs) offer promising opportunities to assist fact-checkers, leveraging LLM’s extensive knowledge and robust reasoning capabilities.
-In this survey paper, we investigate the utilization of generative LLMs in the realm of fact-checking, illustrating various approaches that have been employed and techniques for prompting or fine-tuning LLMs.
-By providing an overview of existing approaches, this survey aims to improve the understanding of utilizing LLMs in fact-checking and to facilitate further progress in LLMs’ involvement in this process.
+The dissemination of false information on online platforms presents a serious societal challenge.
+While manual fact-checking remains crucial, Large Language Models (LLMs) offer promising opportunities to support fact-checkers with their vast knowledge and advanced reasoning capabilities.
+This survey explores the application of generative LLMs in fact-checking, highlighting various approaches and techniques for prompting or fine-tuning these models.
+By providing an overview of existing methods and their limitations, the survey aims to enhance the understanding of how LLMs can be used in fact-checking and to facilitate further progress in their integration into the fact-checking process.
 Generative Large Language Models in Automated Fact-Checking: A Survey
@@
 1 Introduction
-The contemporary digital era has brought various challenges, including the spread of false information.
-The prevalence of social media has made this problem more challenging and a serious societal concern.
-An effective strategy for combating false information involves fact-checking Vlachos and Riedel (2014), mostly done manually by fact-checkers.
-However, the number of fact-checkers is limited, and the spread of false information is increasing Aïmeur et al.
+The modern digital era has introduced numerous challenges, including the widespread dissemination of false information, a problem exacerbated by the rise of social media.
+Fact-checking is a key strategy for combating false information Vlachos and Riedel (2014), but it is still largely carried out manually by fact-checkers.
+As the number of fact-checkers remains insufficient to keep pace with the growing volume of misinformation Aïmeur et al.
+(2023), efforts are underway to develop automated fact-checking systems.
+These systems leverage artificial intelligence and LLMs to reduce the work of fact-checkers Nakov et al.
+(2021a).
+The application of LLMs in automated fact-checking has demonstrated significant potential for improving both the efficiency and accuracy of fact-checking Wang et al.
+(2023a).
+LLMs are trained on large-scale datasets, and they leverage billions of parameters to capture nuances and patterns of natural language.
+Moreover, generative LLMs, specifically designed for text generation, open new possibilities for integrating LLMs into the verification process by generating meaningful, natural text that was not possible before and automating complex reasoning tasks.
+Previous surveys on fact-checking in NLP domain have primarily focused on the needs of fact-checkers Nakov et al.
+(2021a); Hrckova et al.
+(2024), specific fact-checking tasks Guo et al.
+(2022b), datasets Guo et al.
+(2022a) or traditional and BERT-based methods Thorne and Vlachos (2018); Zeng et al.
 (2021).
-Hence, there is an effort to develop an automated fact-checking process by employing digital tools and LLMs that would assist fact-checkers Nakov et al.
-(2021a).
-LLMs utilize large-scale datasets, paired with billions of parameters, to model language nuances and patterns of natural language.
-Furthermore, generative LLMs represent a specialized subset of LLMs tailored for text generation.
-Their application in manual fact-checking has demonstrated promising results in enhancing the efficiency and accuracy of fact-checking.
-With this study, we provide the first comprehensive summarization of approaches and techniques that incorporate generative LLMs in the fact-checking process.
-Existing surveys have explored the needs of fact-checkers and tasks that can be automated Nakov et al.
-(2021a), task definitions, traditional methods, or BERT-like architectures Thorne and Vlachos (2018); Zeng et al.
-(2021).
-Moreover, several authors have examined existing datasets for fact-checking Guo et al.
-(2022).
-Another survey aimed to combat false information using LLMs and explore opportunities and challenges that LLMs bring Chen and Shu (2023b).
-However, this study did not explore methods used in detail, which presents an opportunity to address the role of generative LLMs in fact-checking more closely.
-Our main contribution is to provide an overview of the utilization of generative LLMs in automated fact-checking.
-We present an overview of 70 papers with relevant approaches and novel prompting techniques for researchers who want to delve into LLM-aided information verification.
-We identify four main tasks and various proposed strategies to tackle them.
-Furthermore, we discuss future challenges and possible directions for utilizing LLMs in the information verification process.
+One recent survey addresses combating misinformation using LLMs, illustrating both the opportunities and challenges they present Chen and Shu (2023b).
+However, this study lacks a detailed exploration of the methods employed.
+This presents a gap for a more thorough examination of generative LLMs’ role in fact-checking.
+Our main contribution is a comprehensive overview of approaches and limitations in using generative LLMs for automated fact-checking.
+We review 69 papers, highlighting relevant methodologies and innovative prompting techniques for researchers exploring LLM-aided information verification.
+We analyze fact-checking tasks, LLM methods, common techniques and languages covered in surveyed papers.
+Additionally, we outline future challenges and potential directions for advancing the use of generative LLMs in the verification process.
 2 Methodology
 2.1 Selection
-To identify a set of relevant articles, we employed a manual search of papers from NLP conferences, including ACL, EACL, NAACL and EMNLP.
-Along with these known conferences, we utilized a keyword search using Google Scholar and ArXiv.
-In addition, we extended the created set to include papers referenced in related works and those that cited any article from our retrieved list of relevant papers.
-Additionally, papers submitted to CheckThat!
-Shared tasks were also involved in our survey.
-Overall, we collected 70 papers that employ generative LLMs for fact-checking.
+To compile a set of relevant articles, we conducted a manual search of papers presented at ACL, EACL, NAACL and EMNLP conferences.
+Additionally, we performed a keyword search using Google Scholar and ArXiv.
+We expanded this collection by incorporating papers cited in related works and those that cited any article from our initial list.
+Furthermore, we included papers from the CheckThat!
+Shared tasks.
+In total, we gathered 69 papers that employ generative LLMs for fact-checking.
+Additional details about our selection process are provided in Appendix A.
+2.2 Observed Aspects
+In our survey, we categorized the papers based on several key aspects.
+First, we investigated fact-checking tasks addressed in each study (Section 3) and the methods employed to incorporate LLMs for tackling these tasks (Section 4).
+These methods represent a high-level abstraction of the approaches used, categorized based on the output type generated by LLMs.
+Another aspect we examined is techniques (Section 5), which refer to the strategies, such as prompting and fine-tuning, used to adapt and apply LLMs for specific fact-checking tasks to ensure they generate desired outputs.
+Given that false information is a global issue, we also analyzed languages covered in each study (Section 7).
+Furthermore, we classified what type of knowledge base was used for evidence retrieval (see Table 5) and whether LLMs utilized evidence for the fact-verification task (see Table 5).
+Relevant studies for each task are summarized in Tables 3, 4, 5 and 5.
 2.3 Scope of the Survey
-The survey is intended for researchers with expertise in the domain and a fundamental understanding of methods used across the NLP field.
-Our objective is to present an overview of techniques employed for fact-checking using generative LLMs.
-While fact-checking has evolved to include various modalities, our study delves only into the domain of textual and tabular information.
-Unlike other surveys that describe fact-checking datasets in detail, our focus lies on approaches, techniques, and associated challenges leveraging generative LLMs.
-3 Tasks and Approaches Employed in Fact-Checking
-In this section, we delve into the particular tasks of automated fact-checking, describing its definitions and the methods employed when engaging generative LLMs in the fact-checking process.
-3.1 Check-Worthy Claim Detection
-The task of check-worthy claim detection concerns identifying claims that contain potentially misleading information and that should thus be checked and verified.
-Selecting those claims is challenging, and there is no clear definition of a check-worthy claim.
-However, claims that contain information that may be false or misunderstood and are potentially harmful to society are usually considered to be check-worthy claims.
-Generally, this task is considered a binary classification, the most common approach when employing LLMs.
-However, a finer-grained classification can be used to determine the harmfulness of a claim Wang et al.
+This survey is intended for researchers with expertise in NLP and a fundamental understanding of methods across the field.
+We aim to present an overview of techniques employed in fact-checking using generative LLMs.
+While fact-checking has evolved to cover multiple modalities, our study is confined to textual and tabular data analysis.
+Unlike surveys that provide detailed descriptions of fact-checking datasets, our focus is on the methods, techniques, and challenges of leveraging generative LLMs.
+Our aim is to focus on the verification of the truthfulness of a given input using LLMs, rather than on evaluating LLM outputs for hallucinations Ji et al.
+(2023) or factuality Wang et al.
+(2023a, 2024).
+3 Fact-Checking Tasks
+Fact-checking is the process of verifying the truthfulness of a given claim Guo et al.
 (2022b).
-Additionally, several authors have investigated the generative capabilities of LLMs to generate claims, where LLMs generate a set of verifiable claims from a given text, or questions, which aims to produce questions for further verification Chen et al.
+This process involves several distinct tasks, each essential for determining the credibility of the information.
+In this section, we sort identified tasks based on the number of related papers and highlight the most common methods used for each task.
+3.1 Fact Verification & Fake News Detection
+Fact verification and fake news detection are the most prominent tasks when incorporating LLMs into the fact-checking process, with 53 out of 69 surveyed papers aimed at these tasks.
+Fact verification involves assessing the veracity of a given claim or few sentences.
+In contrast, fake news detection aims to check the trustworthiness of longer texts, such as news articles Li and Zhou (2020).
+[…]
+3.2 Evidence Retrieval
+Thirteen research papers have addressed the task of evidence retrieval using LLMs.
+Evidence retrieval is about gathering essential information from trusted sources to assess the veracity of claims.
+The evidence can take various forms, such as text or tables.
+This evidence serves as the knowledge necessary for accurate fact-checking.
+The selection of an appropriate knowledge source is crucial, as Wikipedia and the open web exhibit several weaknesses.
+In contrast, scientific reports Bhatia et al.
+(2021) offer a more credible alternative.
+[…]
+Figure 1: A taxonomy of methods for integrating generative LLMs in fact-checking, illustrating examples of model inputs and outputs.
+The methods employed with generative LLMs are classified based on the output type.
+3.3 Claim Detection
+Eleven of 69 papers addressed the claim detection task, indicating a moderate emphasis.
+Claim detection concerns identifying claims containing verifiable information necessitating verification.
+These claims, often termed check-worthy or verifiable claims, can include statements that may be false, misleading, or potentially harmful to society.
+[…]
+3.4 Previously Fact-Checked Claims Detection
+The task of previously-checked claims detection aims to reduce redundant work for fact-checkers.
+Currently, it is the least explored area using LLMs, with only five identified papers.
+However, it is essential for fact-checkers to ascertain whether a claim has been previously fact-checked, as misleading claims propagate across various sources and languages Pikuliak et al.
+(2023b).
+Thus, the goal of previously fact-checked claims detection is to identify similar claims from a database that have already undergone fact-checking.
+[…]
+4 Taxonomy of Methods
+In this section, we discuss the methods of using LLMs to perform various fact-checking tasks.
+We categorize methods based on their output type.
+This taxonomy is illustrated in Figure 1: (1) Structured output is used when LLMs process samples and generate answers in a predefined structure: a class for a classification task, a number for a regression task, ranking of inputs, etc.
+In this case, LLMs often replace older specialized approaches.
+(2) Unstructured output is used when LLMs process samples and generate texts such as explanations, summaries, etc.
+that are used in the fact-checking process.
+(3) Synthetic data generation is an approach when LLMs generate new samples for our datasets.
+These datasets can then be used to train other (often smaller and specialized) models.
+4.1 Structured Output
+Methods that produce structured output are the most common in fact-checking, appearing in 59 out of 69 papers.
+These methods are advantageous because they allow easier evaluation, offer higher execution efficiency and easier fine-tuning due to the availability of datasets with labeled targets.
+This makes them particularly beneficial for fact-checking tasks.
+In this section, we further categorize these methods into three output types: classification, regression and ranking.
+Classification.
+LLM-based classification utilizes LLMs to predict categorical labels from a predefined set, e.g., veracity labels.
+In this approach, LLMs are tasked to select the most probable class.
+Classification is commonly employed in fact-checking (54 out of 69), as many tasks involve categorizing claims, especially in claim detection (yes/no) or fact verification (e.g., false, half true, true).
+It is also effective in detecting previously fact-checked claims, as an entailment classification between a given statement and previously fact-checked claims Choi and Ferrara (2023, 2024).
+Regression.
+[…]
+Ranking.
+[…]
+[…]
+4.2 Unstructured Output
+Methods with unstructured outputs typically involve generating continuous text, which poses greater challenges in assessment.
+However, these methods provide more detailed responses, often including justifications, which enhance fact-checking by offering insights into the LLM’s reasoning.
+Since generative LLMs produce open-ended outputs rather than selecting from a set of labels, they can address a broader range of fact-checking tasks.
+We categorized these methods based on the output type generated by LLMs into three categories: (1) coherent, meaningful text consisting of multiple sentences; (2) single-sentence outputs; and (3) outputs consisting of only a few keywords or phrases.
+Explanation Generation.
+[…]
+In addition to directly prompting LLMs to explain their decision, summarization can also be used as a form of explanation generation, in tasks like fact-verification Russo et al.
+(2023b).
+However, the limitation of summarization lies in its tendency to omit crucial details from the retrieved evidence, leading to summaries that may not fully address misleading claims, nor be as persuasive as traditional fact-checking articles.
+Claim and Question Generation.
+The goal of claim generation is usually to produce one sentence that captures all key information for various tasks, including claim detection, evidence retrieval or fact-verification.
+The claim generation can also be known as claim normalization Sundriyal et al.
+(2023), where the aim is to extract a single normalized claim from a social media post.
+However, challenges arise when multiple claims are present in the input, complicating the creation of a single normalized claim.
+Therefore, claim generation can also produce a list of claims.
+A part of claim generation is span detection Gangi Reddy et al.
+(2022), which aims to identify the exact boundaries of the claim within the text.
+The advantage of span detection lies in its ability to select the precise wording of the claim.
+Claim generation and question generation share a similar output format – a single sentence.
+However, in question generation, the output specifically consists of a question, mostly used to verify a given claim Chen et al.
 (2022).
-The advantage of using generative abilities, specifically the above techniques is that LLMs are able to extract key information and decompose complex statements into easily verifiable parts or questions.
-3.2 Previously Fact-Checked Claims Detection
+Query Generation.
+Another form of unstructured output is query generation Prieto-Chavana et al.
+(2023), where the output comprises several words or short phrases primarily used for retrieving evidence from external sources, such as search engines.
+This method allows LLMs to extract essential parts of a claim, enhancing retrieval efficiency and minimizing irrelevant results that may arise from using the full claim.
+Next hop prediction Malon (2021) is a specific type of query generation aimed at identifying and generating the title with a sentence of the evidence to be retrieved.
+This is especially important when previously retrieved evidence introduces new entities that require further elaboration.
+4.3 Synthetic Data Generation
+In synthetic data generation, LLMs are employed to create entire datasets or their parts.
+Unlike the generation from Section 4.2, where LLMs are used to generate results, synthetic data generation creates data that are then used for training or fine-tuning.
+This method proves valuable when datasets are either unavailable or limited for specific tasks or languages.
+In fact-checking, it can facilitate the creation of datasets containing disinformation Chen and Shu (2023a); Vykopal et al.
+(2024) or fake-news Jiang et al.
+(2023); Huang and Sun (2023).
+Synthetic data generation can augment datasets by filling gaps in underrepresented categories or increasing variability.
+For instance, LLM can generate debunked claims to extend datasets Singh et al.
+(2023), produce emotional misinformation tweets Russo et al.
+(2023a) or create counterparts for claim-evidence pairs with contrasting veracity and high word overlap Zhang et al.
+(2024a).
+5 Techniques
+In this section, we analyze how different studies applied techniques to guide LLMs in achieving accurate outputs.
+We observed several popular techniques to improve the performance of LLM, which we categorized into three groups: prompting, fine-tuning and augmentation with external knowledge.
+These techniques can overlap, with some studies using combinations, such as linking prompts with external resources.
+Papers were classified by the most prominent technique; for example, if prompting was combined with external knowledge, they were categorized under augmentation with external knowledge.
+5.1 Prompting
+Prompting is a set of techniques used to improve the performance of LLMs by designing the instructions.
+This approach is versatile and particularly effective in scenarios with limited data, such as fact-checking, where datasets are often scarce and labeled data is difficult to obtain.
+However, prompting faces several limitations, especially with complex tasks like in fact-checking.
+For instance, detecting previously fact-checked claims involves ranking, where prompting alone may struggle to produce accurate results Choi and Ferrara (2024).
+Another limitation is the necessity for clear instructions and contextual descriptions to ensure accurate task execution.
+To reduce ambiguities, it is essential to define all relevant terminology within the prompt, e.g., the attributes of check-worthy claims.
+Therefore, prompt engineering is important, and LLMs can assist in prompt creation Cao et al.
+(2023).
+Prompt structure and wording can influence LLM performance.
+Techniques like role specification Li and Zhai (2023) and JSON formatting Ma et al.
+(2023) have been shown to enhance performance.
+Role specification assigns the LLM a defined role (e.g., fact-checker), improving task comprehension, while JSON formatting structures instructions and outputs in a consistent manner (e.g., Check the ${claim}\nAnswer:$).
+Additionally, output sequencing affect performance Vergho et al.
+(2024); Pelrine et al.
+(2023).
+Presenting the explanation first encourages LLM reasoning, leading to more accurate assessments, whereas placing the final verdict first may cause the LLM to focus on justifying the outcome rather than producing a truthful explanation.
+Fact-checking is challenging for prompting LLMs, as their responses can be inaccurate.
+Zeng and Gao (2023) addressed this by using multiple prompt variants, differing by a single word, to represent different relationships to the claim (e.g., true, unclear, or false).
+This method helps ensure consistency in the LLM’s output, as small changes can affect predictions.
+Moreover, LLMs can be prompted to rewrite claims in a neutral style, improving evaluation reliability Wu and Hooi (2023) and resilience to adversarial attacks.
+Prieto-Chavana et al.
+(2023) used prefixes like Search query: to guide LLMs in generating search queries for evidence retrieval.
+Furthermore, by combining query generation with rationale selection, relevant documents are retrieved, and then the most pertinent information is selected.
+Few-Shot Prompting.
 […]
-In previously fact-checked claims detection, LLMs are usually used to rank the already retrieved fact-checks, e.g., retrieved using the similarity of text embeddings.
-One possible application is to classify pairwise relationships between claims and retrieved fact-checks as textual entailment.
-In addition, LLMs can be employed to filter out irrelevant fact-checks that have been retrieved using another ranking approach.
-3.3 Evidence Retrieval
-The objective of evidence retrieval is to gather crucial information from trusted sources to assess the claim’s veracity, which can have multiple forms, e.g., text or tables.
-The purpose of evidence is to provide the knowledge needed for fact-checking to assess the veracity of the claims that are being fact-checked.
-When incorporating LLMs into the fact-checking, the evidence fills knowledge gaps in LLMs, keeping them updated with up-to-date information they lack.
-In our paper, we refer to this setup as open-book setting, which allows LLMs to access external knowledge, such as Wikipedia or open web Schlichtkrull et al.
-(2023).
-However, selecting the knowledge source is crucial, and Wikipedia and the open web have several weaknesses.
-Thus, scientific reports Bhatia et al.
-(2021) are a more credible choice.
-Conversely, some approaches rely only on internal LLM knowledge and are termed closed-book settings.
-The most employed approach for evidence retrieval is ranking.
-Alternative strategies explored the generative abilities of LLMs for query generation, in which LLM generates a search query for web search, and rationale selection, in which the LLM is tasked to select relevant sentences out of the previously retrieved documents.
-3.4 Fact Verification & Fake News Detection
-Fact verification focuses on determining the truthfulness of a claim or text based on the previously retrieved evidence.
-On the other hand, the fake news detection tries to classify a given input without the evidence based solely on the provided input.
-The most prominent approach employing generative LLMs involves classifying a claim into three categories (supports, refutes, or not enough information) based on the given evidence.
-However, it can also be considered a binary or more fine-grained classification problem, e.g., six Buchholz (2023) or seven categories Hoes et al.
-(2023a).
-In addition to classification, explanation generation is a commonly used technique when incorporating generative LLMs for fact verification, where the goal is to explain the LLM’s decision and clarify what is misleading on a given claim.
-4 Taxonomy of Methods
-Figure 1: A taxonomy of methods employed when incorporating generative LLMs in the fact-checking process.
-Generative LLMs have been employed in fact-checking through several approaches, which we divided into three categories: Classification & Regression, Generation, and Synthetic Data Generation.
-These categories differ in the granularity of the LLM’s output.
-In Classification & Regression, LLMs categorize inputs into discrete labels or continuous values.
-In Generation, the output is continuous text based on the input data.
-For Synthetic Data Generation, LLMs generate parts or entire datasets for training, validating and testing fact-checking methods.
-The proposed taxonomy is illustrated in Figure 1 and a list of relevant papers and detailed information is in Tables 2, 3, 4 and 4.
-Classification & Regression and Generation are further categorized into prompting, fine-tuning and augmentation with external knowledge.
-Prompting leverages instructions to guide responses without LLM’s adaption, fine-tuning adapts LLMs to specific tasks via training on target tasks and knowledge augmentation enhances the responses of LLMs by incorporating up-to-date information.
-4.1 Classification & Regression
-In classification, LLMs predict categorical labels, while in regression LLMs are instructed to provide continuous values as predicted scores on various scales (0-1, 0-100, etc.).
-However, regression is often reframed into a soft classification problem due to the absence of datasets with score targets.
-In soft classification, the authors leverage the produced scores to determine the final class using thresholds (commonly 50%).
-The benefit of soft classification is that the threshold for the classes can be optimized based on the identified properties of LLMs.
-Since the objective of most fact-checking tasks is to classify a claim into a certain category, classification is the most commonly used method in fact-checking, especially in check-worthy claim detection (e.g., yes or no) and fact verification (e.g., false, half true, true).
-Nevertheless, its potential use is also in other tasks, such as detecting previously fact-checked claims as an entailment classification between a given statement and previously fact-checked claims.
-In contrast, regression is widely employed in previously fact-checked claims detection and evidence retrieval, where the identified claims or evidence are ranked using a predicted score.
-Furthermore, regression is leveraged to a small extent in fact verification, intending to predict the truthfulness score for a given claim.
-4.1.1 Prompting
-In the era of generative LLMs, prompting has gained prominence, which is a method of instructing the LLM to condition the response to achieve the desired result.
-This method is employed in many NLP tasks, and several authors investigated the capabilities of prompting in the fact-checking domain.
-In classification, prompts with a list of possible categories are commonly used, while LLMs are instructed to choose the most probable label.
-These prompting techniques are particularly beneficial when datasets are limited across various languages or domains, making traditional fine-tuning methods challenging to use.
-This is especially a concern for fact-checking datasets which are limited mostly by their size.
-Zero-Shot Prompting.
-In zero-shot prompting, LLMs are directly prompted with the instruction without providing examples.
-This is particularly beneficial when there is a lack of extensive training data and quality samples to serve as in-context examples.
-However, zero-shot prompting imposes several limitations.
-One of them is that instructions need to be clear and describe the task context to perform it correctly.
-It is crucial to introduce all the terminology within the prompt to mitigate ambiguities, e.g., attributes of check-worthy claims.
-Due to these challenges, prompt engineering is an essential part of experiments that need to be conducted, especially for fact-checking.
-Another limitation is that the instruction wording and style can significantly affect LLM performance.
-To address this issue, various prompt styles have been proposed.
-In fact-checking, especially role specification Li and Zhai (2023) and JSON format Ma et al.
-(2023) have shown promise in enhancing performance.
-Role specification involves instructing LLM to act based on a defined role, such as fact-checker, while in the JSON format, the instruction adheres to JSON structure (e.g., Check the ${claim}\nAnswer:$).
-Role specification aids LLMs in better understanding the task, while the JSON facilitates easier post-processing due to its structured response format.
-Since fact-checking poses a challenging task for zero-shot prompting and provided answers by LLMs may not always be based on facts, Zeng and Gao (2023) utilized multiple prompt variants to assess the veracity label of each claim.
-These prompt variants differ in only one word and represent various relationships between the claim and whether it is true, unclear or false, to avoid ambiguities in the output of the LLM.
-This technique verifies the consistency of LLMs’ responses to a given claim.
-Few-Shot Prompting.
-Zero-shot prompting has one major limitation: the model is not always able to capture the essence of the task based only on the given instruction.
-Therefore, few-shot prompting is often used to enhance LLM performance by providing demonstrations necessary for complex tasks like fact-checking.
-LLMs efficiently learn the nuances of the task during inference utilizing provided labelled examples.
-Applying few-shot prompting is similar to zero-shot, whereas the authors also investigated the number of samples provided to the LLM Cao et al.
-(2023); Zeng and Gao (2023, 2024).
-Chain-of-Thought.
-The Chain-of-Thought (CoT) technique harnesses the power of LLMs to provide logical reasoning steps.
-This reasoning is performed before the actual class is provided, with the LLM exploiting its reasoning capabilities by incorporating intermediate reasoning steps for the final assessment.
-The CoT can be used in zero-shot Hu et al.
-(2023); Chen and Shu (2023a) and few-shot Liu et al.
-(2024a) settings, where authors include examples of the steps that reasoning should involve, e.g., follow-up questions Sawiński et al.
-(2023).
-4.1.2 Fine-Tuning
-Fine-tuning represents the method for adapting LLMs to a specific task or domain.
-It is mainly employed when the LLM does not have the necessary capabilities to address a given task using prompting, especially in cases where the LLM has no expertise in the selected domain or the task is complex.
-However, due to a large number of parameters, fine-tuning is usually only a concern for smaller LLMs.
-For larger LLMs, fine-tuning is often not feasible due to the limited resources, and also, it is not necessary, since these LLMs have improved capabilities for solving a wide range of tasks, even without further fine-tuning.
-In fact-checking, fine-tuning is primarily employed for ranking tasks, e.g., previously fact-checked claims detection and evidence retrieval.
-However, it is also used for other tasks beyond these applications.
-LLM fine-tuning can be done in multiple settings and one of them is LLM fine-tuning on a single task.
-The advantage of such training is the specialization of the LLM to only a specific task and hence the improved ability of the LLM to address it.
-Conversely, the drawback of this approach lies in the fact that the LLM loses the ability to solve other tasks, even if they are from the same domain (e.g., fact-checking), known as catastrophic forgetting McCloskey and Cohen (1989).
-In this context, addressing multiple tasks necessitates fine-tuning a specific LLM for each task.
-One technique for LLM fine-tuning on a single task involves leveraging the logits of the particular token produced by the LLM, such as true token, which can be used for document ranking Pradeep et al.
-(2021).
-We refer to document ranking based on the probability of the true token as classification ranking because the LLM has been trained in a binary classification setting.
-Building upon classification ranking, multiple authors utilized additional steps Jiang et al.
-(2021); Pradeep et al.
-(2021); Fernández-Pichel et al.
-(2022).
-The classification ranking approach can be further employed to rank evidence in evidence retrieval and to extract pertinent sentences from retrieved evidence Pradeep et al.
-(2020).
-Another option for fine-tuning is multitask fine-tuning, involving multiple tasks during training Du et al.
-(2022).
-This fine-tuning method aims to create a more generalizable model while exploiting knowledge transfer between tasks.
-This technique is not intended to improve the performance of a single task, but of all the tasks involved, where each task can positively affect the others.
-The other benefit compared to training on a single task is that the LLM is specialized for multiple tasks simultaneously and hence it is not necessary to fine-tune separate LLMs for each task.
-LLMs can be fine-tuned with target labels leveraging various prompts.
-For fine-tuning encoder-decoder LLMs, authors mainly employed a prompt of the form "Claim: c Evidence: e Target:" Sarrouti et al.
-(2021), while for fine-tuning decoder-only LLMs, the claim alone is often utilized as input Sawiński et al.
-(2023).
-Since current LLMs are usually instruction-tuned, including the instruction along with the statement is common.
-4.1.3 Augmentation with External Knowledge
-Nevertheless, multiple weaknesses persist in prompting techniques, including a frequent lack of access to up-to-date information.
-Up-to-date information is crucial for accurate fact-checking, and its absence can lead to inaccurate responses.
-To counter this issue, the evidence is commonly incorporated into the prompt.
-Since multiple pieces of evidence may be available to verify a statement’s veracity, there are two schemes that combine multiple pieces of evidence Jiang et al.
-(2021).
-The first method verifies the claim sentence by sentence, and aggregates predicted labels for the final prediction.
-In the second approach, all relevant sentences are employed within a single query, surpassing the performance of the first approach.
-Unlike textual information, evidence can also be in the form of tabular data, which is often linearized before providing it to LLMs Zhang et al.
-(2024b).
-In contrast to incorporating retrieved evidence from the evidence retrieval task, it is common to employ methods for fact-checking, where LLMs interact with external tools, such as search engines or external knowledge bases Zhang and Gao (2023); Yao et al.
-(2023); Quelle and Bovet (2023).
-These techniques provide several benefits, the most significant being that the LLM always have access to the most up-to-date information, which is crucial for fact-checking.
-4.2 Generation
-LLMs exhibit impressive generative abilities, aimed at producing meaningful continuous text.
-This generation capability is beneficial for various NLP tasks, including fact-checking.
-Moreover, the outputs in natural languages enable the integration of LLMs into a wide range of applications, streamlining the fact-checking process.
-Additional advantages include the ability of the LLM to handle diverse topics and to adapt to different contexts, which is particularly valuable for fact-checking where there is a high diversity of topics.
-Generation has the potential to be used across all fact-checking tasks, but especially in the fact verification, to explain misleading information within the claim.
-4.2.1 Prompting
-Prompting in generation involves providing the instruction to guide the LLM in producing meaningful coherent text.
-Prompting techniques are commonly used for generation, due to the absence of labelled data for LLM training, which are challenging to obtain.
-Furthermore, LLMs with billions of parameters are capable of addressing even complex tasks that are part of fact-checking without further fine-tuning.
-Zero-Shot Prompting.
-The same issues arise for the use of zero-shot prompting in generation as were identified for classification and regression, specifically that the style and content of the instruction affect the expected output.
-Since analyzed claims may be challenging for fact verification, LLMs can be instructed to reformulate a given claim into a neutral style that is used for the final assessment Wu and Hooi (2023).
-This approach enhances LLMs’ resistance to adversarial attacks.
-In generation, both the input and the format and order of the output affect the LLM’s performance.
-In fact-checking, where the LLM is instructed to provide an explanation for its assessment, whether the explanation is generated before or after the final assessment is crucial.
-If the final assessment is produced before the explanation, the LLM may focus on matching the final score or category, resulting in an inaccurate explanation.
-Conversely, generating the explanation first allows it to reason behind the LLM’s answer, leading to a more accurate and coherent assessment.
-In zero-shot prompting for evidence retrieval, LLMs can be instructed to generate search queries for search engines to retrieve relevant results.
-For instance, Prieto-Chavana et al.
-(2023) utilized prefixes, e.g., Search query:, to guide LLMs in producing queries.
-This approach allows LLMs to extract essential parts of a claim for better retrieval, avoiding the drawbacks of using the entire claim, which can result in irrelevant evidence.
-Another approach for evidence retrieval is rationale selection, where LLMs aim to select relevant parts of the retrieved evidence or their own knowledge, such as selecting relevant sentences Evans et al.
-(2023); Tan et al.
-(2023).
-Simultaneously, these two approaches complement each other, whereby we can obtain evidence from search engines using generated queries and reduce the evidence to only relevant parts using rationale selection.
-Few-Shot Prompting.
-Few-shot prompting is often used for simpler sub-tasks, such as query or claim generation.
-This approach extends to other tasks, e.g., explanation generation in fact verification.
-However, achieving better performance in these tasks requires quality examples.
-By using a few demonstrations, the LLM can capture how to accurately extract verifiable claims from a given input, model the desired structure of the answer, or recognize the features that generated questions should have.
-Moreover, in-context examples are commonly employed to improve the LLM’s understanding of the task.
-Possible methods of using in-context examples in prompting are span detection Gangi Reddy et al.
-(2022), aiming to identify the exact boundaries of the check-worthy claims within the text, and decomposition of input text into sub-claims Kamoi et al.
+In-context examples are particularly useful in methods such as span detection Gangi Reddy et al.
+(2022) or claim generation Kamoi et al.
 (2023b); Li et al.
 (2023a).
-A few demonstrations can assist LLM in comprehending what characteristics a check-worthy claim is supposed to have, and what its variability might be.
-With this approach, it is important to ensure that the examples capture a wide range of characteristics of check-worthy claims, to eliminate the false positive rate.
-The advantage of span detection lies in the fact that the LLM is instructed to select the exact wording of the claim from the input.
-However, when decomposing the input into sub-claims, the LLM may reformulate the claim potentially altering its meaning and truth value.
-Additionally, few-shot prompting is not limited to only check-worthy claim detection, but is also employed for evidence retrieval, for generating search queries Li et al.
-(2023a) or rationale selection Wang et al.
+To reduce false positives, it is crucial that the examples capture a broad range of characteristics typical for the given task.
+Decomposing input text into sub-claims helps break down complex statements into smaller, verifiable components.
+However, during this process, LLM may unintentionally reformulate claims, potentially altering their meaning or veracity labels.
+By providing varied demonstrations, few-shot prompting can aid LLMs in comprehending the variability and defining characteristics of check-worthy claims.
+Few-shot prompting extends beyond claim detection. It is also employed in evidence retrieval, such as generating search queries Li et al. (2023a) or selecting relevant rationales Wang et al. (2023b).
+Chain-of-Thought.
+[…]
+5.2 Fine-Tuning
+[…]
+Single-Task Fine-Tuning
+One approach to fine-tuning involves fine-tuning on a single task Agrestia et al.
+(2022); Petroni et al.
+(2021); Bhatia et al.
+(2021); Hyben et al.
+(2023); Althabiti et al.
+(2023); Russo et al.
+(2023b); Zeng and Zubiaga (2024), which enhances the model’s specialization and performance in that area.
+However, this method may lead to the phenomenon of catastrophic forgetting, where LLMs lose the ability to perform other tasks, even those within the same domain Luo et al.
+(2024).
+Consequently, addressing multiple tasks requires fine-tuning a specific LLM for each task.
+A strategy for single-task fine-tuning involves classification ranking Pradeep et al.
+(2021).
+This approach has been extended in previous works, e.g., by including named entities within the claim Jiang et al.
+(2021), reranking of previously retrieved documents Pradeep et al.
+(2021) or predicting the reliability of the retrieved passage based on a given claim Fernández-Pichel et al.
+(2022).
+Classification ranking can be employed to rank evidence in evidence retrieval and to extract pertinent sentences from retrieved evidence Pradeep et al.
+(2020).
+Multi-Task Fine-Tuning.
+An alternative approach to single-task fine-tuning involves employing multiple tasks during training Du et al.
+(2022).
+This approach aims to create more generalizable model by exploiting knowledge transfer across tasks.
+While the primary goal is not to enhance performance for a single task, each task can positively affect the others.
+Additionally, this approach allows the LLM to be specialized for multiple tasks simultaneously.
+Prompt Style for Fine-Tuning.
+LLMs can be fine-tuned using target labels and various prompts.
+For fine-tuning encoder-decoder LLMs, researchers commonly employ prompts formatted as "Claim: c Evidence: e Target:" Sarrouti et al.
+(2021).
+In the case of decoder-only LLMs, the input often consists only of the claim Sawiński et al.
+(2023).
+Given that contemporary LLMs are frequently instruction-tuned, incorporating instructions along with the statement is a standard practice.
+Notably, decoder-only models generally outperform encoder-decoder models, especially when trained with instructional inputs.
+Use-cases.
+In fact-checking, fine-tuning is mostly utilized for ranking tasks, e.g., previously fact-checked claims detection and evidence retrieval.
+It also extends to other applications, including generating sub-questions Chen et al.
+(2022) and next-hop prediction Malon (2021).
+5.3 Augmentation with External Knowledge
+Since information changes over time and LLMs often lack up-to-date information, previous techniques are frequently combined with external tools and knowledge bases.
+This augmentation can involve using previously retrieved evidence, web engines, or external databases.
+Evidence in Prompt.
+any approaches integrate evidence from evidence retrieval into prompts to fill LLM knowledge gaps.
+We refer to the use of external knowledge as the open-book setting Schlichtkrull et al.
+(2023), while relying only on internal knowledge is closed-book settings.
+Given that multiple pieces of evidence are often available to verify claim’s veracity Jiang et al.
+(2021), two strategies are used: (1) verifying the claim sentence by sentence and aggregating the predicted labels for the final prediction; or (2) including all relevant evidence within a single query.
+The latter method tends to outperform the former, as consolidating all necessary evidence into one prompt leads to more accurate predictions.
+Beyond textual evidence, tabular data can also serve as a source, and it is typically linearized before being fed into LLMs Zhang et al.
+(2024b).
+Using External Tools.
+Besides incorporating retrieved evidence into prompts, LLMs often interact with external tools, such as search engines or knowledge bases Zhang and Gao (2023); Yao et al.
+(2023); Quelle and Bovet (2023).
+This technique provides several advantages, most notably continuous access to up-to-date information, which is essential for accurate fact-checking.
+However, the effectiveness of these methods depends on the careful selection of credible sources.
+Retrieval-Augmented Generation.
+An effective strategy for mitigating hallucinations and addressing outdated information in LLMs for fact-checking applications is Retrieval Augmented Generation (RAG).
+This method combines information retrieval with LLMs, enhancing their factual accuracy by integrating credible external sources, such as scientific databases and verified fact-checking sites Leippold et al.
+(2024).
+Moreover, RAG systems can be fine-tuned by jointly training the retriever and the LLM, where the LLM provides supervisory signals for training the retriever Izacard et al.
+(2023); Zeng and Gao (2024).
+By combining retrieval and generation, RAG enhances the effectiveness of the fact-checking process against false information.
+6 LLM Pipelines
+In fact-checking, LLMs are commonly integrated into complex orchestrated pipelines, where the output of one step feeds into the next.
+Several studies have proposed such pipelines, combining multiple techniques from Section 5 to aid fact-checkers.
+A common approach is to decompose a claim into verifiable sub-claims and predict the veracity of each sub-claim.
+Zhang and Gao (2023) employed LLMs to break down claims by generating a series of questions and answers.
+The LLM can then rely on its internal knowledge or retrieve external information to answer, assessing its confidence before making a prediction.
+Similarly, Wang and Shu (2023) instructed LLMs to define predicates and follow-up questions.
+After generating answers, the LLM utilize a context of predicted questions and their corresponding answers to assess each predicate’s veracity, which is subsequently aggregated into a final prediction and explanation.
+Alternatively, claims can be assessed without decomposition by directly tasking LLMs to generate questions and answers for veracity prediction and to provide an overall assessment of the claim Chakraborty et al.
+(2023).
+The model then decides whether the claim is supported, refuted or lacks sufficient information.
+Another approach is to convert claims into binary yes-or-no questions, which are used to create interpretable logic clauses for debunking misinformation Liu et al.
 (2024a).
-While, the mentioned advantages, disadvantages and attributes of few-shot prompting are relevant for all fact-checking tasks.
-Chain-of-Thought.
-CoT represents reasoning steps to achieve the desired answer.
-Similarly to classification and regression, CoT can be used in a generation without an example or with several demonstrations.
-However, CoT is not widely employed in generation, since the aim of generation is already to provide meaningful text and explanation is often part of it.
-Few-shot CoT can, for example, be used in a check-worthy claim detection task in producing normalized claim based on a given social media post Sundriyal et al.
-(2023).
-However, the drawback of a single normalized claim arises from the potential presence of multiple facts within these posts, necessitating verification to accurately asses their veracity.
-4.2.2 Fine-Tuning
-Fine-tuning is not as widespread in generation as in classification or regression, since obtaining quality data for this purpose is challenging.
-However, fine-tuning can be used if there is a similar task in another domain with more available data to train LLMs.
-For example, summarization can be leveraged as part of explanation generation in fact verification Russo et al.
-(2023).
-The flaw of summarization for fact-checking is that this approach is unable to capture the essential parts of the retrieved evidence that are necessary to explain what is misleading in the claim and the summaries generated are not as convincing as actual fact-checks.
-In addition, fine-tuning was employed in the generation of sub-questions to verify a given claim Chen et al.
-(2022) and for next-hop prediction Malon (2021), which are both necessary for predicting the veracity label correctly.
-On the one hand, the generation of sub-questions aims to verify the truth of a claim it is necessary to ask the proper questions.
-On the other hand, next-hop prediction is crucial, when retired evidence introduces entities requiring further elaboration.
-4.2.3 Augmentation with External Knowledge
-Researchers have explored various methods to mitigate hallucinations and outdated information in fact-checking.
-One effective method is Retrieval Augmented Generation (RAG), which integrates information retrieval with the LLM, leveraging external sources to enhance accuracy.
-This technique proves powerful in combating hallucinations by using credible information extracted from sources like scientific repositories Leippold et al.
-(2024).
-Moreover, RAG can be fine-tuned by jointly training the retriever and LLM, where the LLM provides supervisory signals for training the retriever Izacard et al.
-(2023); Zeng and Gao (2024).
-4.3 Synthetic Data Generation
-Synthetic data generation aims to generate an entire dataset or its part with LLMs.
-Unlike generation for specific tasks, synthetic data generation precedes the fact-checking pipeline and is not directly involved in fact-checking tasks.
-This approach is advantageous when there are no datasets for some tasks or languages or if there is a limited number of such data.
-In fact-checking, it can be used for generating datasets with disinformation Chen and Shu (2023a); Vykopal et al.
-(2024) or fake-news Jiang et al.
-(2023); Huang and Sun (2023).
-Additionally, synthetic data generation can augment datasets by filling gaps in sparse categories or enhancing variability in the data.
-For instance, LLM can generate debunked claims to extend datasets Singh et al.
-(2023) or create counterparts for claim-evidence pairs with contrasting veracity but high word overlap Zhang et al.
-(2024a).
-5 LLM Pipelines
-It is common that LLMs are used in more complex orchestrated pipelines where the input is being processed in multiple steps and usually the output of one step is used as the input for another.
-Several authors proposed pipelines consisting of multiple fact-checking tasks and employed multiple described methods to provide complex answers necessary for fact-checkers.
-One common approach used by multiple authors consists of decomposing the given claim into easily verifiable sub-claims and predicting the veracity of each generated sub-claim.
-Zhang and Gao (2023) employed the LLM to verify each sub-claim using a generated set of questions and answers.
-In this case, the LLM can utilize its internal knowledge or retrieve information from the web to answer these questions, assessing its confidence level before responding.
-Additionally, Wang and Shu (2023) instructed the LLM to define predicates and follow-up questions.
-Once the answers are provided, the LLM utilizes a context comprising predicted follow-up questions and their answers to assess the veracity of each predicate, which are further used for the final prediction and explanation.
-In addition, the decomposition of the claim into sub-claims can be omitted and the LLM can be directly instructed to generate questions and answers for veracity prediction.
-LLMs can be leveraged to generate questions along with answers, veracity prediction for each answer, and overall veracity label for a given claim Chakraborty et al.
-(2023).
-In this case, the LLM decided whether it supports or refutes the claim or there is not enough information.
-Moreover, claims can be decomposed into binary questions (yes/no) with corresponding answers, while these answers are used as input for generating interpretable logic clauses to debunk misinformation Liu et al.
-(2024a).
-Besides standard generative LLMs, Pan et al.
-(2023b) explored a programming LLM, leveraging its program generation capabilities to create a verification program.
-This approach decomposes the input claim into questions using one generative LLM and then employs programming LLM to generate a reasoning program to guide verification.
-Subsequently, the generated code is executed using the corresponding sub-task function, and the final veracity prediction is produced.
-Since LLMs tend to produce hallucinated and unfaithful explanations, generative LLMs can also be used to enhance faithfulness refinement Kim et al.
+Beyond natural language, Pan et al.
+(2023b) introduced a programming-based LLM for creating verification programs.
+This approach first employs a generative LLM to decompose claims into questions, after which a programming LLM generates a reasoning program for verification.
+The resulting code is executed with specific sub-task functions to produce a final veracity prediction.
+This framework serves as a baseline for fact-verification.
+Since LLMs tend to produce hallucinated and unfaithful explanations, generative LLMs can also be used to enhance faithfulness Kim et al.
 (2024).
@@
 Subsequently, all identified errors and proposed corrections from both debaters are used to adjust the final justification.
-This process is valuable in fact-checking, which requires a convincing and plausible explanation, as a human fact-checker normally would do.
-6 Multilinguality
-False information is not only a problem in English-speaking countries but a worldwide phenomenon, applying LLMs across multiple languages is necessary.
-However, most existing work has focused on applying LLMs to only a single language, especially English.
-In evaluating the performance of LLMs across different languages, most researchers adopted approaches similar to those used in single-language assessment, using the techniques described in Section 4.
-Specifically, when evaluating the fact-checking capability of generative LLMs in various languages, the authors predominantly utilized instructions in a single language, typically English, with the claims presented in their original languages Du et al.
+This process is valuable in fact-checking, which requires convincing and plausible explanations and fact-check articles, as a human fact-checker normally would do.
+7 Languages
+We also examined the language coverage of each work.
+Most papers focus on a single language, primarily English ($56\times$56 ×), with only one each for Chinese and Arabic.
+Only 11 papers addressed more than one language, typically pairing English with another.
+Notably, only three explored ten or more languages, with the largest experiments covering 114 languages Setty (2024).
+Prompt language.
+Most researchers use approaches designed for single-language settings in multilingual contexts, typically providing instructions in English, while leaving the claims in their original languages Du et al.
 (2022); Agrestia et al.
@@
 (2023); Li and Zhai (2023).
-This methodology allowed for the use of the same prompts without adapting it to the language of the analyzed claim.
-Moreover, the instruction can be in the same language as the analyzed claim, which has been observed to enhance the accuracy of LLMs in identifying false information Hu et al.
-(2023).
-However, this approach has its drawbacks.
-It requires predefined instructions for each supported language, and it also relies on accurate language detectors.
-Another method employed in the case of multilingualism and the use of generative LLMs, focuses on fine-tuning LLMs on multilingual data, to create a generalized LLM capable of processing inputs in multiple languages Quelle and Bovet (2023).
-Additionally, LLMs can be trained in one language, particularly a high-resource one, and then evaluated in other languages, a process known as cross-lingual transfer Pikuliak et al.
-(2021).
+This allows for consistent prompts without adaption.
+Alternatively, instructions can be in the claim’s language, which requires predefined instructions for each language and an accurate language detector.
+Multilingual training.
+An alternative approach involves fine-tuning LLMs on multilingual data to create models capable of handling multiple languages Quelle and Bovet (2023).
+Another possible strategy is cross-lingual transfer, where models are trained on high-resource languages and then evaluated in other languages.
 8 Future Directions and Challenges
 Knowledge-Augmented Strategies.
-Techniques like ReAct Yao et al.
-(2023) and RAG Lewis et al.
-(2021), which incorporate external knowledge or tools, have shown promise in addressing complex NLP tasks, including fact-checking.
-Despite their potential, these methods are still underexplored in fact-checking, only 5 out of 70 surveyed papers.
-Future directions lie in harnessing these techniques to enhance the efficiency and accuracy of fact-checking, which brings opportunities for detecting false information.
-Multilingual Fact-Checking Benchmarks.
-Evaluating LLMs in a multilingual context is necessary for fact-checking since false information is a global issue.
-Developing fact-checking benchmarks in multiple languages can help to assess and improve the LLM’s cross-lingual capabilities.
-Since new LLMs emerge rapidly, the established benchmarks will aid in faster and more efficient evaluation of both current and future LLMs, thus empowering fact-checkers to always use state-of-the-art models.
+Techniques incorporating external knowledge into LLMs have shown promise in addressing complex tasks in the NLP domain, including fact-checking.
+Despite their potential, these techniques remain underexplored in fact-checking, with only five out of 69 surveyed papers employing external sources.
+Future research should focus on leveraging these techniques, e.g., RAG, to enhance the efficiency and accuracy of LLM-driven fact-checking.
+Multilingual Fact-Checking.
+Given the global nature of false information, leveraging LLMs for fact-checking across languages is essential.
+Developing effective multilingual fact-checking techniques will enhance the ability of LLMs to detect and address misinformation in diverse linguistic contexts.
+One challenge includes the underexplored task of previously fact-checked claims detection, which also involves cross-lingual claim-matching – where input claims and fact-checks are in different languages.
+This presents a valuable opportunity for further research, highlighting the potential of LLMs in multilingual fact-checking.
 Fact-Checking in Real Time.
-In the current research landscape, most approaches act reactively, often after a claim gains traction on social media platforms.
-Leveraging generative LLMs can enhance this process by enabling continuous monitoring and instant analysis of emerging information across various platforms.
-To ensure LLMs keep up-to-date and effective in fact-checking, they need to be integrated with tools allowing LLMs to access the newest information in real-time.
-This proactive approach can identify and flag false claims and enhance the accuracy and reliability of information dissemination.
+Current methods are mostly reactive, addressing claims after they gain traction on social media.
+Generative LLMs could enable real-time monitoring and analysis, providing continuous updates on emerging information.
+Integrating LLMs with real-time information will help to identify and flag false information as it arises, enhancing the timeliness and accuracy of false claim detection and mitigating misinformation more effectively.
 Interactive Fact-Checking.
-Interactive fact-checking tools, especially those exploiting generative LLMs, represent a promising future direction in combating false information.
-Users can engage in dialogue with an LLM to explore the accuracy of claims, ask follow-up questions or request explanations of its decisions.
-The development of interactive fact-checking tools offers an excellent opportunity to empower fact-checkers in effective false information mitigation.
+Interactive fact-checking tools powered by generative LLMs represent a promising direction for users to verify claims through dynamic dialogues.
+These systems could facilitate deeper engagement with fact-checking, offering explanations and follow-up questions.
+Future work should explore the development of interactive fact-checking tools to empower fact-checkers in false information mitigation.
 9 Conclusion
-The rapid development of generative LLMs has stimulated considerable interest in their potential utilization in fact-checking.
-Through a systematic review of 70 research papers, this study gave an overview of methods that incorporate LLMs into information verification.
-This survey discusses the techniques and approaches explored in the reviewed literature, which not only provides a comprehensive overview of current methodologies but also presents the basis for future research efforts to improve existing methods and explore new frontiers in LLM-assisted fact-checking.
+The rapid advancement of generative LLMs has sparked considerable interest in their potential applications within the fact-checking domain.
+This study presents a systematic review of 69 research papers, providing a detailed analysis of the various methods that incorporate generative LLMs into the information verification process.
+By examining the techniques and approaches explored in the existing literature, this survey offers a comprehensive overview of current methodologies and a foundation for future research efforts to enhance and explore new frontiers in LLM-assisted fact-checking.
 Limitations
 Focus on Generative LLMs.
-This survey paper aims to provide an overview of the approaches and techniques employed in fact-checking while covering only generative LLMs.
-There are various types of language models, but in recent years, the majority of the research in the NLP domain has focused on LLMs, especially on generative LLMs.
+This survey paper focuses exclusively on approaches and techniques employed in fact-checking that leverage generative LLMs.
+While various types of language models exist, such as encoder-only models, recent research in NLP has increasingly centered on generative LLMs due to their rapid advancements and versatility across a wide range of tasks.
+Although other model architectures could contribute to fact-checking, this paper emphasizes generative LLMs due to their growing prominence and potential in this domain.
+LLMs for Fact-Checking.
+Using LLMs in fact-checking intersects with several related domains, including hallucination, LLM factuality or truth discovery.
+These domains address the issue of false information from various perspectives.
+However, most research on hallucination and LLM factuality aims to evaluate the model’s outputs and determine whether there are discrepancies between generated answers and real-world facts.
+In contrast, this survey focuses on approaches and techniques for verifying information rather than evaluating LLM outputs.
+Additionally, we include methods for improving the faithfulness of LLM-generated explanations to enhance the reliability of fact-checking.
 Textual Information Only.
-Although false information is not only an issue in textual form, but it can involve various modalities, such as images, videos, speech, etc., in our survey we focus only on the problem of verifying textual information.
-Therefore, we excluded all papers that did not involve text processing in their approach.
+False information extends beyond text and can involve various modalities, such as images, videos, and audio.
+However, this survey is limited to the problem of verifying textual information.
+We excluded studies focused on multimodal approaches, thereby narrowing our scope to fact-checking methods that involve text processing only.
 Acknowledgements
@@
 APVV-21-0114.
-Appendix A Datasets & Languages
+Appendix A Methodology
+To search for research papers, we employed mostly search on ACL Anthology [^1] https://github.com/acl-org/acl-anthology/, ArXiv and Google Scholar.
+Since the fact-checking domain is still evolving and there are multiple names for the same fact-checking tasks, we used task names from previous surveys on fact-checking.
+Based on fact-checking tasks, we defined a list of keywords for the search that are shown in Table 1.
+The search for articles was completed in February 2024.
+After filtering out the papers that are not directly related to fact-checking, are focused on fact-checking the LLM’s output or do not use generative LLMs for fact-checking, we identified a list of 69 articles.
+Keywords
+fact-checking
+check-worthy claim detection
+claim detection
+previously fact-checked claims detection
+previously fact-checked claims retrieval
+claim-matching
+evidence retrieval
+factuality prediction
+claim verification
+fact verification
+fake-news detection
+justification prediction
+explanation generation
+fact-checking dataset
+disinformation detection
+misinformation detection
+rumour verification
+rumour detection
+Table 1: A list of keywords used when searching for research papers.
+Appendix B Datasets & Languages
 After the rise of social media, the necessity for the development of datasets aimed at verifying information began to emerge.
@@
 MultiClaim Pikuliak et al.
-(2023b), X-Fact Gupta and Srikumar (2021) or Pinocchio Hu et al.
-(2023).
+(2023b) or X-Fact Gupta and Srikumar (2021).
 A list of all identified datasets employed for the LLM experiments is shown in Table LABEL:tab:datasets.
-Concerning linguistic scope, the vast majority of papers are concentrated on English ($57\times$57 ×) or a single language like Chinese ($1\times$1 ×) or Arabic ($1\times$1 ×).
-Out of the total papers surveyed, 12 analyzed more than one language, predominantly combining English with another language.
-Notable, only four investigated the employing LLMs across ten or more languages.
-Nevertheless, the most extensive language coverage in LLM experiments involved 114 languages Setty (2024), or 46 languages in the fact verification task.
-Appendix B Fact-Checking Benchmarks
+Appendix C Fact-Checking Benchmarks
 Several benchmarks have been developed to assess the performance of LLMs in fact-checking tasks.
-Hu et al.
-(2023) introduced the Pinocchio benchmark, encompassing seven domains related to factual knowledge and reasoning, evaluating ten LLMs with different prompting strategies.
-In contrast, Bang et al.
+For example, Bang et al.
 (2023) focused on assessing ChatGPT’s reasoning, hallucination, and interactivity capabilities across eight tasks, including misinformation detection.
@@
 (2024).
-Appendix C Model Analysis
+Appendix D Model Analysis
 Various researchers have focused on different types of LLMs within the reviewed papers.
-In contrast, these could be divided into two main categories: Encoder-Decoder (Seq2Seq) and Decoder-only (as shown in Figure 2).
+In contrast, these could be divided into two main categories: Encoder-Decoder (Seq2Seq) and Decoder-only.
 Although the first Seq2Seq models already emerged in 2019, their first applications in fact-checking were seen in 2021, when the problem of automated fact-checking began to receive increasing attention.
-The most commonly used Seq2Seq models include the T5 model, while we also encounter the T0, Flan-T5, or multilingual mT5.
-Figure 2: The hierarchy of identified models in surveyed papers.
+The most commonly used Seq2Seq models include the T5 model, while we also encounter the T0 Sanh et al.
+(2022), Flan-T5 Chung et al.
+(2022), or mT5 Xue et al.
+(2021).
 In contrast, the second category includes Decoder-only LLMs, which we further divided into GPT-base, LLaMA-based and Other models, similarly to Minaee et al. (2024).
@@
 This category includes all GPT models (e.g.
-GPT-3, GPT-3.5, or GPT-4), CODEX.
+GPT-3, GPT-3.5, or GPT-4) or CODEX.
 In addition to these closed-source models, there are also several open-sourced models, such as the GPT-Neo model.
@@
 Along with these models, several authors also investigated Atlas, a retrieval-augmented language model.
-Figure 3 depicts LLMs employed in more than five papers, accompanied by the respective count of papers each model utilized. A total of 35 various LLMs were employed in the collection of 70 papers.
+Figure 2 depicts LLMs employed in more than five papers, accompanied by the respective count of papers each model utilized. A total of 33 various LLMs were employed in the collection of 69 papers.
 Figure 2: The occurrence frequency of the models in the reviewed papers. We only present models used in more than five papers.
 Table 2: A list of datasets utilized in experiments with LLMs for fact-checking.
-Name     Check-worthy claim detection     Previously fact-checked claims detection     Evidence retrieval     Fact verification # Lang. Size Citation
+Name     Claim detection     Previously fact-checked claims detection     Evidence retrieval     Fact verification # Lang. Size Citation
 ArCOV19-Rumors 1 138 Haouari et al. (2021)
@@
 BoolQ-FV / AmbiFC 1 11k Glockner et al. (2023)
+ChatGPT-FC 1 22k Li et al. (2023b)
+Check-COVID 1 2k Wang et al. (2023b)
 CHEF 1 10k Hu et al. (2022)
+ClaimBuster 1 24k Arslan et al. (2020)
+ClaimDecomp 1 1k Chen et al. (2022)
 CLAN 1 6k Sundriyal et al. (2023)
@@
 CLIMATE-FEVER 1 2k Diggelmann et al. (2021)
-COVID-19 Scientific 1 142 Lee et al. (2020)
-COVID-Fact 1 4k Saakyan et al. (2021)
-ChatGPT-FC 1 22k Li et al. (2023b)
-Check-COVID 1 2k Wang et al. (2023a)
-ClaimBuster 1 24k Arslan et al. (2020)
-ClaimDecomp 1 1k Chen et al. (2022)
 Climate Feedback 1 N/A -
@@
 Constraint 1 11k Patwa et al. (2021)
+COVID-19 Scientific 1 142 Lee et al. (2020)
+COVID-Fact 1 4k Saakyan et al. (2021)
 Data Common X N/A -
 e-FEVER 1 68k Stammbach and Ash (2020)
-EX-FEVER 1 61k Ma et al. (2024)
 EnvClaims 1 29k Stammbach et al. (2023)
 ExClaim 1 4k Gurrapu et al. (2022)
+EX-FEVER 1 61k Ma et al. (2024)
+FactEX 1 12k Althabiti et al. (2023)
+FakeNewsNet 1 23k Shu et al. (2019)
 FEVER 1 185k Thorne et al. (2018)
@@
 FM2 1 13k Eisenschlos et al. (2021)
-FactEX 1 12k Althabiti et al. (2023)
-FakeNewsNet 1 23k Shu et al. (2019)
 FullFact 1 N/A -
@@
 HoVer 1 26k Jiang et al. (2020)
+Labeled Unreliable News 1 74k Rashkin et al. (2017)
 LESA-2021 1 379k Gupta et al. (2021)
 LIAR 1 13k Wang (2017)
-LIAR++ 1 6k Russo et al. (2023)
+LIAR++ 1 6k Russo et al. (2023b)
 LIAR-NEW 2 2k Pelrine et al. (2023)
 LLMFake 1 100 Chen and Shu (2023a)
-Labeled Unreliable News 1 74k Rashkin et al. (2017)
 Monant Medical Misinformation Dataset 1 51k Srba et al. (2022)
@@
 NewsClaims 1 889 Gangi Reddy et al. (2022)
-Pinocchio 27 21k Hu et al. (2023)
 PolitiFact 1 N/A -
@@
 Symmetric 1 1k Schuster et al. (2019)
+TabFact 1 18k Chen et al. (2020)
 TREC 2020 Health Misinfo. 1 46 Clarke et al. (2020)
 TREC-COVID 1 171k Roberts et al. (2021)
-TabFact 1 18k Chen et al. (2020)
 Twitter15 1 2k Liu et al. (2015)
 Twitter16 1 778 Ma et al. (2016)
+VerMouth 1 12k Russo et al. (2023a)
 VitaminC 1 489k Schuster et al. (2021)
@@
 X-Fact 26 31k Gupta and Srikumar (2021)
-Paper Approaches Techniques # Lang.
+Paper Methods Techniques # Lang.
 Gangi Reddy et al. (2022) - classification - span detection - fine-tuning - zero-shot - few-shot 1
@@
 Du et al. (2022) - binary classification - fine-tuning 6
-Chen et al. (2022) - claim decomposition - fine-tuning 1
+Chen et al. (2022) - question generation - fine-tuning 1
 Sawiński et al. (2023) - binary classification - fine-tuning - zero-shot - few-shot - few-shot CoT 1
-Kamoi et al. (2023b) - text generation - few-shot 1
-Li et al. (2023a) - text generation - few-shot 1
+Kamoi et al. (2023b) - claim generation - few-shot 1
+Li et al. (2023a) - claim generation - few-shot 1
 Hyben et al. (2023) - binary classification - fine-tuning - zero-shot - few-shot 10
 Cao et al. (2023) - binary classification - zero-shot - few-shot - zero-shot CoT 2
-Wang et al. (2023b) - 3-class classification - zero-shot 1
-Sundriyal et al. (2023) - text generation - few-shot - zero-shot CoT 1
+Sundriyal et al. (2023) - claim generation - few-shot - zero-shot CoT 1
 Setty (2024) - binary classification - zero-shot 114
-Table 2: Work addressing the task of check-worthy claim detection in the context of fact-checking. Each paper includes the utilized approaches, techniques and the number of languages analyzed.
-Paper Approaches Techniques # Lang.
+Table 3: Work addressing the task of claim detection in the context of fact-checking. Each paper includes the utilized approaches, techniques and the number of languages analyzed.
+Paper Methods Techniques # Lang.
 Shliselberg et al. (2022) - ranking task - fine-tuning 1
@@
 Table 4: A list of papers addressing the task of detecting previously fact-checked claims. We include approaches, techniques and a number of languages for each paper.
-Paper Approaches Techniques Knowledge base # Lang.
+Paper Methods Techniques Knowledge base # Lang.
 Pradeep et al. (2020) - ranking - rationale selection - fine-tuning - MS-MARCO 1
@@
 Schlichtkrull et al. (2023) - question generation - few-shot - search engine 1
-Wang et al. (2023a) - rationale selection - few-shot - scientific papers 1
+Wang et al. (2023b) - rationale selection - few-shot - scientific papers 1
 Evans et al. (2023) - rationale selection - zero-shot - scientific papers 1
@@
 {longtabu}
-to lllll Paper Approaches Techniques Used Evidence # Lang.
+to lllll Paper Methods Techniques Used Evidence # Lang.
 Pradeep et al. (2020) - 3-class classification - fine-tuning - open-book 1
@@
 Buchholz (2023) - 6-class classification - fine-tuning - zero-shot - closed-book 1
-Wang et al. (2023a) - 3-class classification - few-shot - open-book 1
+Wang et al. (2023b) - 3-class classification - few-shot - open-book 1
 Pan et al.
@@
 Guan et al. (2023) - regression - fine-tuning - zero-shot - open-book - closed-book 1
-Hu et al. (2023) - 3-class classification - zero-shot - few-shot - zero-shot CoT - few-shot CoT - self-refinement - self-consistency - open-book 27
 Wu and Hooi (2023) - binary classification - zero-shot - closed-book 1
@@
 Pan et al. (2023a) - binary classification - explanation generation - zero-shot - few-shot - few-shot CoT - open-book - closed-book 1
-Russo et al. (2023) - explanation generation - fine-tuning - open-book 1
+Russo et al. (2023a) - data generation - zero-shot 1
+Russo et al. (2023b) - explanation generation - fine-tuning - open-book 1
 Li and Zhai (2023) - 6-class classification - zero-shot - closed-book 2